### v1

# Language Models Encode Collaborative Signals in Recommendation

Abstract
Recent studies empirically indicate that language models (LMs) encode rich world knowledge beyond mere semantics, attracting significant attention across various fields. However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information. Contrary to the prevailing understanding that LMs and traditional recommender models learn two distinct representation spaces due to a huge gap in language and behavior modeling objectives, this work rethinks such understanding and explores extracting a recommendation space directly from the language representation space. Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance. This outcome suggests the homomorphism between the language representation space and an effective recommendation space, implying that collaborative signals may indeed be encoded within advanced LMs. Motivated by these findings, we propose a simple yet effective collaborative filtering (CF) model named AlphaRec, which utilizes language representations of item textual metadata (e.g., titles) instead of traditional ID-based embeddings. Specifically, AlphaRec is comprised of three main components: a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function, making it extremely easy to implement and train. Our empirical results show that AlphaRec outperforms leading ID-based CF models on multiple datasets, marking the first instance of such a recommender with text embeddings achieving this level of performance. Moreover, AlphaRec introduces a new language-representation-based CF paradigm with several desirable advantages: being easy to implement, lightweight, rapid convergence, superior zero-shot recommendation abilities in new domains, and being aware of user intention. Codes are available at https://github.com/LehengTHU/AlphaRec.
1 Introduction
(a) Linearly mapping language representations into the recommendation space
(b) Performance comparison
(c) The t-SNE representations of movies and user intention in two spaces.
Figure 1: Linearly mapping item titles in language representation space into recommendation space yields superior recommendation performance on Movies & TV [1] dataset. (1(a)) The framework of linear mapping. (1(b)) The recommendation performance comparison between leading CF recommenders and linear mapping. (1(c)) The t-SNE [2] visualizations of movie representations, with colored lines linking identical movies or user intention across language space (left) and linearly projected recommendation space (right).
Language models (LMs) have achieved great success across various domains [3, 4, 5, 6, 7], prompting a critical question about the knowledge encoded within their representation spaces. Recent studies empirically find that LMs extend beyond semantic understanding to encode comprehensive world knowledge about various domains, including game states [8], lexical attributes [9], and even concepts of space and time [10] through language modeling. However, in the domain of recommendation where the integration of LMs is attracting widespread interest [11, 12, 13, 14, 15], it remains unclear whether LMs inherently encode relevant information on user preferences and behaviors. One possible reason is the significant difference between the objectives of language modeling for LMs and user behavior modeling for recommenders [16, 17, 18, 19].
Currently, one prevailing understanding holds that general LMs and traditional recommenders encode two distinct representation spaces: the language space and the recommendation space (i.e., user and item representation space), each offering potential enhancements to the other for recommendation tasks [17, 20]. On the one hand, when using LMs as recommenders, aligning the language space with the recommendation space could significantly improve the performance of LM-based recommendation [14, 21, 22, 23]. Various alignment strategies are proposed, including fine-tuning LMs with recommendation data [15, 16, 24, 25, 26], incorporating embeddings from traditional recommenders as a new modality of LMs [17, 20, 27], and extending the vocabulary of LMs with item tokens [18, 19, 28, 29, 30, 31]. On the other hand, when using LMs as the enhancer, traditional recommenders greatly benefit from from leveraging text representations [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], semantic and reasoning information [46, 47, 48, 49], and generated user behaviors [50, 51]. Despite these efforts, explicit explorations of the relationship between language and recommendation spaces remain largely unexplored.
In this work, we rethink the prevailing understanding and explore whether LMs inherently encode user preferences through language modeling. Specifically, we test the possibility of directly deriving a recommendation space from the language representation space, assessing whether the representations of item textual metadata (e.g., titles) obtained from LMs can independently achieve satisfactory recommendation performance. Positive results would imply that user behavioral patterns, such as collaborative signals (i.e., user preference similarities between items) [52, 53], may be implicitly encoded by LMs. To test this hypothesis, we employ linear mapping to project the language representations of item titles into a recommendation space (see Figure 1(a)). Our observations include:
  • Surprisingly, this simple linear mapping yields high-quality item representations, which achieve exceptional recommendation performance (see Figure 1(b) and experimental results in Section 2).
  • The clustering of items is generally preserved from the language space to the recommendation space (see Figure 1(c)). For example, movies with the theme of superheroes and monsters are gathering in both language and recommendation spaces.
  • Interestingly, the linear mapping effectively reveals preference similarities that may be implicit or even obscure in the language space. For instance, while certain movies, such as those of homosexual movies (illustrated in Figure 1(c)), show dispersed representations in the language space, their projections through linear mapping tend to cluster together, reflecting their genres affiliation.
These findings indicate the homomorphism [54] between the language representation space of LMs and an effective item representation space for recommendation. Motivated by this insight, we propose a new language-representation-based recommendation paradigm for general collaborative filtering (CF), which utilizes the pre-trained language representations of item titles as the item input and the average historical interactions’ representations as the user input. Different from traditional ID-based CF models [55, 56, 52] that heavily rely on trainable user and item IDs, this paradigm solely uses pre-trained LM embeddings and completely abandons ID-based embeddings. In this paper, to fully explore the potential of advanced language representations, we adopt a simple model architecture consisting of a two-layer MLP with graph convolution, and the popular contrastive loss, InfoNCE [57, 58, 59], as the objective function. This model is named AlphaRec for its originality, a series of good properties, and an adjustable hyperparameter $\alpha$ for user intention capture.
Benefiting from paradigm shifts from ID-based embeddings to language representations, AlphaRec presents three desirable advantages. First, AlphaRec is notable for its simplicity, lightweight, rapid convergence, and exceptional recommendation performance (see Section 4.1). We empirically demonstrate that, for the first time, such a simple model with embeddings from pre-trained LMs can outperform leading CF models on multiple datasets. This finding strongly supports the possibility for developing language-representation-based recommender systems. Second, AlphaRec exhibits a strong zero-shot recommendation capability across untrained datasets (see Section 4.2). By co-training on three Amazon datasets (Books, Movies & TV, and Video Games) [1], AlphaRec can achieve performance comparable to the fully-trained LightGCN on entirely different platforms (MovieLens-1M [60] and BookCrossing [61]), and even exceed LightGCN in a completely new domain (Amazon Industrial), without additional training on these target datasets. This capability underscores AlphaRec’s potential to develop more general recommenders. Third, AlphaRec is user-friendly, offering a new research paradigm that enhances recommendation by leveraging language-based user feedback (see Section 4.3). Endowed with its inherent semantic comprehension of language representations, AlphaRec can refine recommendations based on user intentions expressed in natural language, enabling traditional CF recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
2 Uncovering Collaborative Signals in LMs via Linear Mapping
In this section, we aim to explore whether LMs implicitly encode user preference similarities in their representation spaces. We first formulate the personalized item recommendation task, then detail the linear mapping method and its empirical findings. Empirical evidence indicates the homomorphism [54] between the representation spaces of advanced LMs and effective recommendation spaces.
Task formulation. Personalized item recommendation with implicit feedback aims to select items $i\in\mathcal{I}$ that best match user $u$’s preferences based on binary interaction data $\mathbf{Y}=[y_{ui}]$, where $y_{ui}=1$ ($y_{ui}=0$) indicates user $u\in\mathcal{U}$ has (has not) interacted with item $i$ [59]. The primary objective of recommendation is to model the user-item interaction matrix $\mathbf{Y}$ using a scoring function $\hat{y}:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$, where $\hat{y}_{ui}$ measures $u$’s preference for $i$. The scoring function $\hat{y}_{ui}=s\circ\phi_{\theta}(\mathbf{x}_{u},\mathbf{x}_{i})$ comprises three key components: pre-existing features $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$ for user $u$ and item $i$, a representation learning module $\phi_{\theta}(\cdot,\cdot)$ parametrized by $\theta$, and a similarity function $s(\cdot,\cdot)$. The representation learning module $\phi_{\theta}$ transfers $u$ and $i$ into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ for similarity matching $s(\cdot,\cdot)$, and the Top-$K$ highest scoring items are recommended to $u$.
Different recommenders employ various pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$ and representation learning architecture $\phi_{\theta}(\cdot,\cdot)$. Traditional ID-based recommenders use one-hot vectors as pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$. The choice of ID-based representation learning architecture $\phi_{\theta}$ can vary widely, including ID-based embedding matrix [55], multilayer perception [62], graph neural network [52, 63], and variational autoencoder [64]. We adopt a commonly used similarity function, cosine similarity [65, 58] $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, in this paper.
Linear mapping. Linear mapping ensures the homomorphism [54] between the vector spaces, and is therefore widely used in the study on the representation space properties of LMs [10, 66, 67]. In this section, we adopt a trainable linear mapping matrix $\bm{W}$ as the representation learning module $\phi_{\theta}$, mapping representations from the language space into an item recommendation space. High performance on the test set would indict the homomorphism between the language space and an effective recommendation space, which further suggests the possible existence of collaborative signals (i.e., user preference similarities between items) [56] in the language representation space [68, 10]. The overall framework of the linear mapping process is illustrated in Figure 1(a). Specifically, we utilize LMs as feature extractors, leveraging the language representations of item titles as initial item feature $\mathbf{x}_{i}$. For initial user feature $\mathbf{x}_{u}$, we use the average of the title representations of historically interacted items, defined as $\mathbf{x}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{x}_{i}$, where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with. Detailed procedures for obtaining these language-based features are provided in Appendix B.2. The linear mapping matrix sets user and item representations as $\mathbf{e}_{u}=\bm{W}\mathbf{x}_{u}$ and $\mathbf{e}_{i}=\bm{W}\mathbf{x}_{i}$ respectively. To learn the linear mapping matrix $\bm{W}$, we adopt the InfoNCE loss [57] as the objective function, which has demonstrated state-of-the-art performance in both ID-based [69, 70] and LM-enhanced CF recommenders [47] (refer to Equation (4) for the formula). We directly use linearly mapped representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ to calculate the user-item similarity $s(\mathbf{e}_{u},\mathbf{e}_{i})$ for recommendation.
Empirical findings. We compare the recommendation performance of the linear mapping method with three classical CF baselines, matrix factorization (MF) [55, 71], MultVAE [64], and LightGCN [56] (see more details about baselines in Appendix C.2.1). We report three widely used metrics Hit Ratio (HR@$K$), Recall@$K$, Normalized Discounted Cumulative Gain (NDCG@$K$)) to evaluate the effectiveness of linear mapping, with $K$ set by default at 20. We evaluate a wide range of LMs, including BERT-style models [4, 5], decoder-only large language models (LLMs) [6, 72], and LLM-based text embedding models [73, 74] (see Appendix B.1 for details about used LMs).
Table 1: The recommendation performance of linear mapping comparing with classical CF baselines.
Books Movies & TV Video Games
Recall NDCG HR Recall NDCG HR Recall NDCG HR
CF MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $\underline{0.3418}$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
LightGCN (He et al., 2021) $0.0723$ 0.0608 0.3489 $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
Linear Mapping BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
RoBERTa 0.0247 0.0209 0.1262 0.0406 0.0387 0.2277 0.0578 0.0338 0.1339
Llama2-7B 0.0662 0.0559 0.3176 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746
Mistral-7B 0.0650 0.0544 0.3124 0.1039 0.0963 0.4994 0.1270 0.0687 0.2428
text-embedding-ada-v2 0.0515 0.0436 0.2570 0.0926 0.0874 0.4563 0.1176 0.0683 0.2579
text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
SFR-Embedding-Mistral 0.0738 0.0610 0.3371 0.1152 0.1065 0.5327 0.1370 0.0787 0.2927
Table 1 reports the recommendation performance yielded by the linear mapping on three Amazon datasets [1], comparing with classic CF baselines. We observe that the performance of most advanced text embedding models (e.g., text-embeddings-3-large [73] and SFR-Embedding-Mistral [74]) exceeds LightGCN on all datasets. We also empirically show that these improvements may not merely come from the better feature encoding ability (refer to Appendix B.3). Figure 1(c) further suggests that obscure user preference similarities may be implicitly encoded in the language representation space, with a linear mapping matrix to activate. These findings indicate the homomorphism [54] between the language representation spaces of advanced LMs and effective item representation spaces for recommendation. In addition, with the help of linear mapping, for the first time, we are able to clearly quantify the gap in the representation spaces between different LMs for recommendation. With the advances in LMs, we observe that the performance of language representations exhibits a rising trend, especially showing a sharp rise from BERT-style models to LLMs. Representations from early BERT-style models (e.g., BERT [4] and RoBERTa [5]) only show weaker or equal capabilities compared with MF, failing behind LLMs largely. This further suggests that, until very recently, the language representation spaces are able to encode user preference similarities efficiently.
3 AlphaRec
This finding of space homomorphism [54] sheds light on building advanced CF models purely based on language representations without introducing ID-based embeddings. Inspired by this, we develop a simple yet effective CF model called AlphaRec, by incorporating only three simple components (i.e., nonlinear projection [62], graph convolution [56] and contrastive learning (CL) objectives [57]). It is important to highlight that our approach is centered on exploring the potential of language representations for recommendation by integrating essential components from leading CF models, rather than deliberately inventing new CF mechanisms. Surprisingly, these basic and simple components lead to an excellent language-representation-based recommender with multiple capabilities. We present the model structure of AlphaRec in Section 3.1, and compare AlphaRec with two popular recommendation paradigms in Section 3.2.
3.1 Method
We present how AlphaRec is designed and trained. Generally, the representation learning architecture $\phi_{\theta}(\cdot,\cdot)$ of AlphaRec is simple, which only contains a two-layer MLP and the basic graph convolution operation, with language representations as the input features $\mathbf{x}_{u},\mathbf{x}_{i}$. The cosine similarity is used as the similarity function $s(\cdot,\cdot)$, and the contrastive loss InfoNCE [57, 58] is adopted for optimization. For simplicity, we consistently adopt text-embeddings-3-large [73] as the language representation model, for its excellent language understanding and representation capabilities.
Nonlinear projection. In AlphaRec, we substitute the linear mapping matrix delineated in Section 2 with a nonlinear MLP. This conversion from linear to nonlinear is non-trivial, for the paradigm shift from ID-based embeddings to language representations, since nonlinear transformation helps in excavating more comprehensive user preference similarities from the language representation space with rich semantics (see discussions about this in Appendix C.2.3) [62]. Specifically, we project the language representation $\mathbf{x}_{i}$ of the item title to an item space for recommendation with the two-layer MLP, and obtain user representations as the average of historical items:
$\mathbf{e}_{i}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{2},\quad\mathbf{e}_{u}^{(0)}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{e}^{(0)}_{i}.$ (1)
Graph convolution. Graph neural networks (GNNs) have shown superior effectiveness for recommendation [52, 56], owing to the natural user-item graph structure in recommender systems [75]. In AlphaRec, we employ a minimal graph convolution operation [56] to capture more complicated collaborative signals from high-order connectivity [56, 76, 77, 75] as follows:
$\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\mathbf{e}_{i}^{(k)},\quad\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\mathbf{e}_{u}^{(k)}.$ (2)
The information of connected neighbors is aggregated with a symmetric normalization term $\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}$. Here $\mathcal{N}_{u}$ ($\mathcal{N}_{i}$) denotes the historical item (user) set that user $u$ (item $i$) has interacted with. The features $\mathbf{e}^{(0)}_{u}$ and $\mathbf{e}^{(0)}_{i}$ projected from the MLP are used as the input of the first layer. After propagating for $K$ layers, the final representation of a user (item) is obtained as the average of features from each layer:
$\mathbf{e}_{u}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{u}^{(k)},\quad\mathbf{e}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{i}^{(k)}.$ (3)
Contrastive learning objective. The introduction of contrasting learning is another key element for the success of leading CF models. Recent research suggests that the contrast learning objective, rather than data augmentation, plays a more significant role in improving recommendation performance [70, 78, 69]. Therefore, we simply use the contrast learning object InfoNCE [57] as the loss function without any additional data augmentation on the graph [79, 58]. With cosine similarity as the similarity function $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, the InfoNCE loss [57, 79, 80] is written as:
$\mathbf{\mathop{\mathcal{L}}}_{\text{InfoNCE}}=-\sum_{(u,i)\in\mathcal{O}^{+}}\log\frac{\exp{(s(\mathbf{e}_{u},\mathbf{e}_{i})/\tau)}}{\exp{(s(\mathbf{e}_{u},\mathbf{e}_{i})/\tau)}+\sum_{j\in\mathcal{S}_{u}}\exp{(s(\mathbf{e}_{u},\mathbf{e}_{j})/\tau)}}.$ (4)
Here, $\tau$ is a hyperparameter called temperature [81], $\mathcal{O}^{+}=\{(u,i)|y_{ui}=1\}$ denoting the observed interactions between users $\mathcal{U}$ and items $\mathcal{I}$. And $\mathcal{S}_{u}$ is a randomly sampled subset of negative items that user $u$ does not adopt.
3.2 Discussion of Recommendation Paradigms
We compare the language-representation-based AlphaRec with two popular recommendation paradigms in Table 2 (see more discussion about related works in Appendix A).
ID-based recommendation (ID-Rec) [52, 55]. In the traditional ID-based recommendation paradigm, users and items are represented by ID-based learnable embeddings derived from a large number of user interactions. While ID-Rec exhibits excellent recommendation capabilities with low training and inference costs [63, 79], it also has two significant drawbacks. Firstly, these ID-based embeddings learned in specific domains are difficult to transfer to new domains without overlapping users and items [37], thereby hindering zero-shot recommendation capabilities. Additionally, there is a substantial gap between ID-Rec and natural languages [34], which makes ID-based recommenders hard to incorporate language-based user intentions and further refine recommendations accordingly.
LM-based recommendation (LM-Rec) [15, 16, 24]. Benefitting from the extensive world knowledge and powerful reasoning capabilities of LMs [7, 82], the LM-based recommendation paradigm has gained widespread attention [11, 13]. LM-Rec tends to convert user interaction history into text prompts as input for LMs, utilizing pre-trained or fine-tuned LMs in a text generation pattern to recommend items. LM-Rec demonstrates zero-shot and few-shot abilities and can easily understand language-based user intentions. However, LM-Rec faces significant challenges. Firstly, the LM-based model architecture leads to huge training and inference costs, with real-world deployment difficulties. Additionally, limited by the text generation paradigm, LM-based models tend to perform candidate selection [17] or generate a single next item [24]. It remains difficult for LM-Rec to comprehensively rank the entire item corpus or recommend multiple items that align with user interests.
Language-representation-based recommendation. We argue that AlphaRec follows a new CF paradigm, which we term the language-representation-based paradigm. This paradigm replaces the ID-based embeddings in ID-Rec with representations from pre-trained LMs, employing feature encoders to map language representations directly into the recommendation space. Few early studies lie in this paradigm, including using BERT-style LMs to learn universal sequence representations [37, 44], or adopting the same model architecture as ID-Rec with simple input features replacement [34, 35]. These early explorations, which are mostly based on BERT-style LMs, are usually only applicable in certain specific scenarios, such as the transductive setting with the help of ID-based embeddings [37]. This phenomenon is consistent with our previous findings in Section 2, indicating that BERT-style LMs may fail to effectively encode user preference similarities. We point out that AlphaRec is the first recommender in the language-representation-based paradigm to surpass the traditional ID-based paradigm on multiple tasks, faithfully demonstrating the effectiveness and potential of this paradigm.
Table 2: Comparison of recommendation paradigms
Recommendation Paradigms Training Cost Zero-shot Ability Intention-aware Ability
ID-based Low
LM-based High
Language-representation-based Low
4 Experiments
In this section, we aim to explore the effectiveness of AlphaRec. Specifically, we are trying to answer the following research questions:
  • RQ1: How does AlphaRec perform compared with leading ID-based CF methods?
  • RQ2: Can AlphaRec learn general item representations, and achieve good zero-shot recommendation performance on entirely new datasets?
  • RQ3: Can AlphaRec capture user intention described in natural language and adjust the recommendation results accordingly?
4.1 General Recommendation Performance (RQ1)
Motivation. We aim to explore whether the language-representation-based recommendation paradigm can outperform the ID-Rec paradigm. An excellent performance of AlphaRec would shed light on the research line of building language-representation-based recommenders in the future.
Baselines. We consider ID-based baselines in this section. LM-based methods are not incorporated for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection or next item prediction. In addition to classic baselines (i.e., MF, MultVAE, and LightGCN) introduced in section 2, we consider two categories of leading ID-based CF baselines: CL-based CF methods: SGL [83], BC Loss [79], XSimGCL [70] and LM-enhanced CF methods: KAR [48], RLMRec [47]. See more details about baselines in Appendix C.2.1.
Results. Table 3 presents the performance of AlphaRec compared with leading CF baselines. The best-performing methods are bold, while the second-best methods are underlined. Figure 2(a) and Figure 2(b) report the training efficiency and ablation results. We observe that:
Table 3: The performance comparison with ID-based CF baselines. The improvement achieved by AlphaRec is significant ($p$-value <<much-less-than$<$ < 0.05).
Books Movies & TV Video Games
Recall NDCG HR Recall NDCG HR Recall NDCG HR
MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $0.3418$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
LightGCN (He et al., 2021) $0.0723$ $0.0608$ $0.3489$ $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
SGL (Wu et al., 2021) 0.0789 0.0657 0.3734 0.0916 0.0838 0.4680 0.1089 0.0634 0.2449
BC Loss (Zhang et al., 2022) 0.0915 0.0779 0.4045 0.1039 0.0943 0.5037 0.1145 0.0668 0.2561
XSimGCL (Yu et al., 2024) 0.0879 0.0745 0.3918 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550
KAR (Xi et al., 2023) 0.0852 0.0734 0.3834 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571
RLMRec (Ren et al., 2024a) 0.0928 0.0774 0.4092 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997
AlphaRec 0.0991* 0.0828* 0.4185* 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207*
Imp.over the best baseline $6.79\%$ $5.34\%$ $2.27\%$ $9.12\%$ $10.75\%$ $5.40\%$ $9.75\%$ $10.51\%$ $7.01\%$
(a) Training efficiency comparison
(b) Ablation study on Books
Figure 2: (2(a)) The bar charts show the number of epochs needed for each model to converge. AlphaRec tends to exhibit an extremely fast convergence speed. (2(b)) The effect of each component in AlphaRec on Books dataset.
  • AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets. AlphaRec shows an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline RLMRec [47]. We further conduct the ablation study to explore the reason for its success (see more ablation results in Appendix C.2.2). As shown in Figure 2(b), each component in AlphaRec contributes positively. Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively. Moreover, the performance drop from replacing InfoNCE loss [58] with BPR loss [71] (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance. These findings suggest that, by carefully designing the model to extract user preference similarities, the potential of language-representation-based paradigm can be realized, surpassing the leading ID-Rec paradigm.
  • The incorporation of semantic language representations into traditional ID-based CF methods can lead to significant performance improvements. We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over CL-based CF methods. Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec. We attribute this phenomenon to the fact that the performance contribution of these methods mainly comes from the language representations, which is consistent with the previous findings [34, 44].
  • AlphaRec exhibits fast convergence speed. We find that the convergence speed of AlphaRec is comparable with, or even surpasses, CL-based methods with data augmentation (e.g., SGL [83] and XSimGCL [70]). Meanwhile, methods based solely on graph convolution (LightGCN [56]) or CL objective (BC Loss [79]) show relatively slow convergence speed, indicating that introducing these modules may not lead to convergence speed improvement. We attribute the fast convergence speed of AlphaRec to the homomorphism between the language space and a good recommendation space. As a result, only minor adjustments to the vanilla language representations are needed for effective recommendation.
4.2 Zero-shot Recommendation Performance on Entirely New Datasets (RQ2)
Motivation. We aim to explore whether AlphaRec has learned general item representations [37], which enables it to perform well on entirely new datasets without any user and item overlap.
Task and datasets. In zero-shot recommendation [38], there is not any item or user overlap between the training set and test set [38, 33], which is different from the research line of cross-domain recommendation in ID-Rec [84]. We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M [60], Book Crossing [61], and Industrial [1]) without further training on these new datasets. (see more details about how we train AlphaRec on multiple datasets in Appendix C.3.1).
Baselines. Due to the lack of zero-shot recommenders in the field of general recommendation, we slightly modify two zero-shot methods in the sequential recommendation [85], ZESRec [37] and UniSRec [37], as baselines. We also incorporate two strategy-based CF methods, Random and Pop (see more details about these baselines in Appendix C.3.2).
Table 4: The zero-shot recommendation performance comparison on entirely new datasets. The improvement achieved by AlphaRec is significant ($p$-value <<much-less-than$<$ < 0.05).
Industrial MovieLens-1M Book Crossing
Recall NDCG HR Recall NDCG HR Recall NDCG HR
full MF (Rendle et al., 2012) 0.0344 0.0225 0.0521 0.1855 0.3765 0.9634 0.0316 0.0317 0.2382
MultVAE (Liang et al., 2018) 0.0751 0.0459 0.1125 0.2039 0.3741 0.9740 0.0736 0.0634 0.3716
LightGCN (He et al., 2021) 0.0785 0.0533 0.1078 0.2019 0.4017 0.9715 0.0630 0.0588 0.3475
zero-shot Random 0.0148 0.0061 0.0248 0.0068 0.0185 0.2611 0.0039 0.0036 0.0443
Pop 0.0216 0.0087 0.0396 0.0253 0.0679 0.5439 0.0119 0.0101 0.1157
ZESRec (Ding et al., 2021) 0.0326 0.0272 0.0628 0.0274 0.0787 0.5786 0.0155 0.0143 0.1347
UniSRec (Hou et al., 2022) 0.0453 0.0350 0.0863 0.0578 0.1412 0.7135 0.0396 0.0332 0.2454
AlphaRec 0.0913* 0.0573 0.1277* 0.1486* 0.3215* 0.9296* 0.0660* 0.0545* 0.3381*
Imp.over the best zero-shot baseline $157.09\%$ $127.69\%$ $30.29\%$ $66.67\%$ $64.16\%$ $37.78\%$ $101.55\%$ $63.71\%$ $47.97\%$
Results. Table 4 presents the zero-shot recommendation performance comparison on entirely new datasets. The best-performing methods are bold and starred, while the second-best methods are underlined. We observe that:
  • AlphaRec demonstrates strong zero-shot recommendation capabilities, comparable to or even surpassing the fully trained LightGCN. On datasets from completely different platforms (e.g., MovieLens-1M and Book Crossing), AlphaRec is comparable with the fully trained LightGCN. On the same Amazon platform dataset, Industrial, AlphaRec even surpasses LightGCN, for which we attribute to the possibility that AlphaRec implicitly learns unique user behavioral patterns on the Amazon platform [1]. Conversely, ZESRec and UniSRec exhibit a marked performance decrement compared with AlphaRec. We attribute this phenomenon to two aspects. On the one hand, BERT-style LMs [4, 5] used in these works may not have effectively encoded user preference similarities, which is consistent with our previous findings in Section 2. On the other hand, components designed for the next item prediction task in sequential recommendation [86] may not be suitable for capturing the general preferences of users in CF scenarios.
  • The zero-shot recommendation capability of AlphaRec generally benefits from an increased amount of training data, without harming the performance on source datasets. As illustrated in Figure 8, the zero-shot performance of AlphaRec, when trained on a mixed dataset, is generally superior to training on one single dataset [37]. Additionally, we also note that training data with themes similar to the target domain contributes more to the zero-shot performance. For instance, the zero-shot capability on MovieLens-1M may primarily stem from Movies & TV. Furthermore, we discover that AlphaRec, when trained jointly on multiple datasets, hardly experiences a performance decline on each source dataset. These findings further point to the general recommendation capability of a single pre-trained AlphaRec across multiple datasets. The above findings also offer a potential research path to achieve general recommendation capabilities, by incorporating more training data with more themes. See more details about these results in Appendix C.3.3.
4.3 User Intention Capture Performance (RQ3)
Motivation. We aim to investigate whether a straightforward paradigm shift enables pre-trained AlphaRec to perceive text-based user intentions and refine recommendations. Specifically, we introduce an adjustable hyperparameter $\alpha$ to combine user intentions with historical interests.
Task and datasets. We test the user intention capture ability of AlphaRec on MovieLens-1M and Video Games. In the test set, only one target item remains for each user [87], with one intention query generated by ChatGPT [88, 40] (see the details about how to generate and check these intention queries in Appendix C.4.1). In the training stage, we follow the same procedure as illustrated in Section 2 to train AlphaRec. In the inference stage, we obtain the language representation $\mathbf{e}_{u}^{Intention}$ for each user intention query and combine it with the original user representation to get a new user representation as $\mathbf{\tilde{e}}_{u}^{(0)}=(1-\alpha)\mathbf{e}_{u}^{(0)}+\alpha\mathbf{e}_{u}^{Intention}$ [87]. This new user representation is sent into the freezed AlphaRec for recommendation. We report a relatively small $K$ = 5 for all metrics to better reflect the intention capture accuracy.
Table 5: The performance comparison in user intention capture.
MovieLens-1M Video Games
HR@5 NDCG@5 HR@5 NDCG@5
TEM (Bi et al., 2020) 0.2738 0.1973 0.2212 0.1425
AlphaRec (w/o Intention) 0.0793 0.0498 0.0663 0.0438
AlphaRec (w Intention) 0.4704* 0.3738* 0.2569* 0.1862*
(a) Case study of user intention capture
(b) Effect of $\alpha$
Figure 3: User intention capture experiments on MovieLens-1M. (3(a)) AlphaRec refines the recommendations according to language-based user intention. (3(b)) The effect of user intention strength $\alpha$.
User intention capture results. Table 5 represents the user intention capture experiment results, compared with the baseline TEM [89]. Clearly, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention). Moreover, AlphaRec outperforms the baseline model TEM by a large margin, even without additional training on search tasks. We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user (see more case study results in Appendix C.4.3). As shown in Figure 3(a), AlphaRec accurately captures the hidden user intention for “Godfather”, while keeping most of the recommendation results unchanged. This indicates that AlphaRec captures the user intention and historical interests simultaneously.
Effect of the intention strength $\alpha$. By controlling the value of $\alpha$, AlphaRec can provide better recommendation results, with a balance between user historical interests and user intent capture. Figure 3(b) depicts the effect of $\alpha$. Initially, as $\alpha$ increases, the recommendation performance rises accordingly, indicating that incorporating user intention enables AlphaRec to provide better recommendation results. However, as the $\alpha$ approaches 1, the recommendation performance starts to decrease, which suggests that the user historical interests learned by AlphaRec also play a vital role. The similar effect of $\alpha$ on Video Games is discussed in Appendix C.4.4.
5 Limitations
There are several limitations not addressed in this paper. On the one hand, although we have demonstrated the excellence of AlphaRec for multiple tasks on various offline datasets, the effectiveness of online employment remains unclear. On the other hand, although we have successfully explored the potential of language-representation-based recommenders by incorporating essential components in leading CF models, we do not elaboratively focus on designing new components for CF models.
6 Conclusion
In this paper, we explored what knowledge about recommendations has been encoded in the LM representation space. Specifically, we found that the advanced LMs representation space exhibits a homomorphic relationship with an effective recommendation space. Based on this finding, we developed a simple yet effective CF model called AlphaRec, which exhibits good recommendation performance with zero-shot recommendation and user intent capture ability. We pointed out that AlphaRec follows a new recommendation paradigm, language-representation-based recommendation, which uses language representations from LMs to represent users and items and completely abandons ID-based embeddings. We believed that AlphaRec is an important stepping stone towards building general recommenders in the future. [^1] Thc broader impact of AlphaRec will be detailed in Appendix E
Appendix A Related Works
Representations in LMs. The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space. An important and effective approach for interpreting and analyzing representations of LMs is linear probing [68, 67]. The main idea of linear probing is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure [9] ) from the representations in the hidden layers of LMs. A high probing result (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations [68, 10]. Recent studies empirically demonstrate that concepts such as color [90], game states [8]. and geographic position are encoded in LMs. Furthermore, these concepts may even be linearly encoded in the representation space of LMs [8, 91].
Collaborative filtering. Collaborative filtering (CF) [92] is an advanced technique in modern recommender systems. The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup [55]. Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern [71]. Recent advances in CF mainly benefit from two aspects, graph convolution [75] and contrastive learning [92]. These CF models exhibit superior recommendation performance by conducting the embedding propagation [52, 56] and applying contrastive learning objectives [83, 63, 70]. However, although effective, these methods are still limited, due to the ID-based paradigm. Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains [37] or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) [34].
LMs for recommendation. The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation. The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LLM-based recommendation. The first research direction, LLM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs [48, 47, 46, 49, 93, 94]. Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences. The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) [34, 35]. These methods tend to train the LM as the text modality encoder together with the traditional recommender. In previous studies, BERT-style LMs are widely used as the text modality encoder. The third research line, LLM-based recommendation, directly uses LLMs as the recommender and recommends items in a text generation paradigm. Early attempts focus on adopting in-context learning (ICL) [95] and prompting pre-trained LLMs [96, 97, 98, 99]. However, such naive methods tend to yield poor performance compared to traditional models. Therefore, recent studies concentrate on fine-tuning LLMs on recommendation-related corpus [16, 15, 26, 25, 29] and align the LLMs with the representations from traditional recommenders as the additional modality [17, 20, 27, 100].
Appendix B Linear Mapping
B.1 Brief of Used LMs
We briefly introduce the LMs we use for linear mapping in Section 2.
  • BERT [4] is an encoder-only language model based on the transformer architecture [3], pre-trained on text corpus with unsupervised tasks. BERT adopts bidirectional self-attention heads to learn bidirectional representations.
  • RoBERTa [5] is an enhanced version of BERT. RoBERTa preserves the architecture of BERT but improves it by training with more data and large batches, adopting dynamic masking, and removing the next sentence prediction objective.
  • Llama2-7B [6] is an open-source decoder-only LLM with 7 billion parameters. Llama2 adopts grouped-query attention, with longer context length and larger size of the pre-training corpus compared with Llama-7B [101].
  • Mistral-7B [72] is an open-source pre-trained decoder-only LLM with 7 billion parameters. Mistral 7B leverages grouped-query attention, coupled with sliding window attention for faster and lower cost inference.
  • text-embedding-ada-v2 & text-embeddings-3-large [73] are leading text embedding models released by OpenAI. These models are built upon decoder-only GPT models, pre-trained on unsupervised data at scale.
  • SFR-Embedding-Mistral [74] is a decoder-based text embedding model built upon the open-source LLM Mixtral-7B [72]. SFR-Embedding-Mistral introduces task-homogeneous batching and computes contrastive loss on “hard negatives”, which brings a better performance than the vanilla Mixtral-7B model.
B.2 Extracting Representations from LMs
We present how to extract representations from LMs. For encoder-based LMs (e.g., BERT [4] and RoBERTa [5]), we use the representation of the last hidden state corresponding to the [CLS] token [40]. For decoder-based models (e.g., Llama-7B [6, 72], Mistral-7B, and SFR-Embedding-Mistral [74]), we use the representation in the last transformer block [3], corresponding to the last input token [10, 102, 73]. Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^2] https://platform.openai.com/docs/guides/embeddings [73]), we directly call the API interface to obtain representations.
B.3 Empirical Findings
We examinate whether the better performance of leading LMs comes from the more compact representation ability. We randomly shuffle item representations and conduct the same linear mapping experiment. As illustrated in Table 6, randomly shuffled representations, text-embeddings-3-large (Random), lags largely behind the vanilla linear mapping method, suggesting that the performance improvement not merely comes from the better feature encoding ability.
Table 6: Linear mapping performance of randomly shuffled item representations
Books Movies & TV Video Games
Recall NDCG HR Recall NDCG HR Recall NDCG HR
BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
text-embeddings-3-large (Random) 0.0200 0.0197 0.1316 0.0559 0.0528 0.3204 0.0562 0.0328 0.1351
text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
Appendix C Experiments
C.1 Datasets
Table 7: Dataset statistics.
Books Movies & TV Video Games Industrial MovieLens-1M Book Crossing
#Users 71,306 26,073 40,834 15,141 6,040 6,273
#Items 26,073 12,464 14,344 5,163 3,043 5,335
#Interactions 2,209,030 876,027 390,013 82,578 995,492 253,057
Density 0.0008 0.0026 0.0007 0.0010 0.0542 0.0076
We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com [1] (i.e., Books, Movies & TV, Video Games, and Industrial), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing). Table 7 reports the data statistics of each dataset.
We divide the history interaction of each user into training, validation, and testing sets with a ratio of 4:3:3, and remove users with less than 20 interactions following previous studies [50]. We also remove items from the test and validation sets that do not appear in the training set, to address the cold start problem.
In this paper, we only use the item titles as the text description. Figure 4 gives some item title examples from different datasets.
Item Title Examples Books: Dismissed with Prejudice: A J.P. Beaumont Novel; Die for Love: A Jacqueline Kirby Novel of Suspense; The Cloud; Memories Before and After the Sound of Music: An Autobiography; Harry Potter and the Sorcerer’s Stone; Movies & TV: Batman Begins; Fantastic Four; Max Headroom: The Complete Series; Madagascar; Land of the Dead; King Kong; Video Games: Fighting Force; Tomb Raider II; Tomb Raider; WWF Warzone; Kartia: The Word of Fate; Snowboard Kids; Command & amp; Conquer: Tiberian Sun - PC; Final Fantasy VII; Grim Fandango - PC; Half-Life - PC; MovieLens-1M: Basquiat (1996); Tin Cup (1996); Godfather, The (1972); Supercop (1992); Manny & Lo (1996); Bound (1996); Carpool (1996); Book Crossing: Prague : A Novel; Chocolate Jesus; Wie Barney es sieht; To Kill a Mockingbird; Sturmzeit. Roman; A Soldier of the Great War; Pride and Prejudice (Dover Thrift Editions); Industrial: Jurassic Perisphinctes Ammonites from France; FS9140: Spinosaurus - Dinosaur Tooth 20-30mm; FS9410: USA Eocene, Fossil Fish (Knightia alt), A-grade; Delta 50-857 Charcoal Filter for 50-868; Hitachi RP30SA 7-1/2 Gallon Stainless Steel Industrial Shop Vacuum (Discontinued by Manufacturer); Makita 632002-4 14-Inch Cut-Off Wheels (5-Pack) (Discontinued by Manufacturer); PORTER-CABLE 740001801 4 1/2-Inch by 10yd 180 Grit Adhesive-Backed Sanding Roll;
Figure 4: Example of item titles.
C.2 General Recommendation
C.2.1 Baselines
We incorporate a series of CF models as our baselines for general recommendation. These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec). For these LM-enhanced CF methods, we adopt the leading CF method XSimGCL as the backbone.
  • MF [55, 71] is the most basic CF model. It denotes users and items with ID-based embeddings and conducts matrix factorization with Bayesian personalized ranking (BPR) loss.
  • MultVAE [64] is a traditional CF model based on the variational autoencoder (VAE). It regards the item recommendation as a generative process from a multinomial distribution and uses variational inference to estimate parameters. We adopt the same model structure as suggested in the paper: 600 → 200 → 600.
  • LightGCN [56] is a light graph convolution network tailored for the recommendation, which deletes redundant feature transformation and activation function in NGCF [52].
  • SGL [83] introduces graph contrastive learning into recommender models for the first time. By employing node or edge dropout to generate augmented graph views and conduct contrastive learning between two views, SGL achieves better performance than LightGCN.
  • BC Loss [79] introduces a robust and model-agnostic contrastive loss, handling various data biases in recommendation, especially for popularity bias.
  • XSimGCL [70] directly generates augmented views by adding noise into the inner layer of LightGCN without graph augmentation. The simplicity of XSimGCL leads to a faster convergence speed and better performance.
  • KAR [48] enhances recommender models by integrating knowledge from large language models (LLMs). It generates textual descriptions of users and items and combine the LM representations with traditional recommenders using a hybrid-expert adaptor.
  • RLMRec [47] aligns semantic representations of users and items with the representations in CF models through a contrastive loss, as an additional loss trained together with the CF model. The fusion of semantic information and collaborative information brings performance improvement.
C.2.2 Ablation Study
(a) Ablation study on Movies & TV
(b) Ablation study on Video Games
Figure 5: Ablation study
We conduct the same ablation study as introduced in Section 4.1 on Movies & TV and Video Games datasets. As illustrated in Figure 5, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.1.
C.2.3 The t-SNE Visualization Comparison
(a) LM representations
(b) AlphaRec (w/o MLP)
(c) AlphaRec
Figure 6: The t-SNE visualization of representations on Movies & TV. (6(a)) The item representations in the LM space. (6(b)) The item representations obtained by replacing the MLP with a linear mapping matrix in AlphaRec. (6(c)) The item representations obtained from AlphaRec.
In this section, we aim to intuitively explore how the MLP in AlphaRec further helps in excavating collaborative signals in language representations, compared to the linear mapping matrix. We visualize the item representations from LMs, AlphaRec (w/o MLP), and AlphaRec in Figure 6, where AlphaRec (w/o MLP) denotes replacing the MLP with a linear mapping matrix. We observed that movies about superhero and monster cluster in all representation spaces, indicating both AlphaRec (w/o MLP) and AlphaRec capture the preference similarities between these items and preserve the clustering relationship. The difference between AlphaRec (w/o MLP) and AlphaRec lies in the ability to capture obscure preference similarities among items. As shown in Figure 6(a), homosexual movies are dispersed in the language space, indicating the possible semantic differences between them. AlphaRec successfully captures the preference similarities and gathers these items in the representation space, while AlphaRec (w/o MLP) remains some items dispersed. Moreover, AlphaRec outperforms AlphaRec (w/o MLP) by a large margin, as indicated in Figure 5(a). These results indicate that AlphaRec exhibits a more fine-grained preference capture ability with the help of nonlinear transformation.
C.3 Zero-shot Recommendation
C.3.1 Co-training on Multiple Datasets
Co-training on multiple datasets is similar to training on one single dataset, where the only difference lies in the negative sampling. When co-training on multiple datasets, the negative items are restricted to the same dataset as the positive item rather than the full item pool. The other training procedures remain the same with training on one single dataset.
C.3.2 Baselines
Since previous works about zero-shot recommendation mostly focus on sequential recommendation [86, 85], we slightly modify two methods in sequential recommendation, ZESRec [38] and UniSRec [37] as our baselines. Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
  • Random denotes randomly recommending items from the entire item pool.
  • Pop denotes randomly recommending from the most popular items. Here popularity denotes the number of users that have interacted with the item.
  • ZESRec [38] is the first work that defines the problem of zero-shot recommendation. To address this problem, this work introduces a hierarchical Bayesian model with representations from the pre-trained BERT.
  • UniSRec [37] aims to learn universal item representations from BERT, with parametric whitening and a MoE-enhanced adaptor. By pre-training on multiple source datasets, UniSRec can conduct zero-shot recommendation on various datasets in a transductive or inductive paradigm.
C.3.3 The Effect of Training Datasets
The effect of the training dataset scale on zero-shot recommendation. We report the zero-shot recommendation performance differences trained on different datasets in Table 8. Here AlphaRec (trained on Books) denotes training on a single Books dataset, while AlphaRec (trained on mixed dataset) denotes co-training on three Amazon datasets. Generally, training on more datasets leads to a better zero-shot performance. In addition, we observe that, for the zero-shot performance on untrained target datasets, training datasets with similar themes contribute more (e.g., Movies & TV and MovieLens-1M).
The performance comparison between training on the single dataset and the mixed dataset. In Table 9, AlphaRec (trained on single dataset) denotes training and testing on the same single dataset, while AlphaRec (trained on mixed dataset) denotes training on three Amazon datasets (i.e., Books, Movies & TV, and Video Games) and testing on one single dataset. Generally, co-training on three Amazon datasets yields similar performance compared with training on one single dataset. The only exception lies in Video Games, which shows some performance degradation. We attribute this to the difference between the selection of $\tau$. We use $\tau$ = 0.15 when trained on the mixed dataset, while the optimal $\tau$ for Video Games lies around 0.2. These results indicate that a single AlphaRec can capture user preferences among various datasets, showcasing a general collaborative signal capture ability.
Table 8: The effect of the training dataset on zero-shot recommendation
Industrial MovieLens-1M Book Crossing
Recall NDCG HR Recall NDCG HR Recall NDCG HR
AlphaRec (trained on Books) 0.0896 0.0562 0.1256 0.1218 0.2619 0.8942 0.0646 0.0532 0.3346
AlphaRec (trained on Movies & TV) 0.0909 0.0581 0.1266 0.1438 0.3122 0.9200 0.0471 0.0406 0.2600
AlphaRec (trained on Video Games) 0.0905 0.0567 0.1225 0.1221 0.2313 0.9034 0.0412 0.0378 0.2585
AlphaRec (trained on mixed dataset) 0.0913 0.0573 0.1277 0.1486 0.3215 0.9296 0.0660 0.0545 0.3381
Table 9: Performance comparison between training on the single dataset and the mixed dataset
Books Movies & TV Video Games
Recall NDCG HR Recall NDCG HR Recall NDCG HR
AlphaRec (trained on single dataset) 0.0991 0.0828 0.4185 0.1221 0.1144 0.5587 0.1519 0.0894 0.3207
AlphaRec (trained on mixed dataset) 0.0979 0.0818 0.4147 0.1194 0.1107 0.5463 0.1381 0.0827 0.2985
C.4 User Intention Capture
C.4.1 Intention Query Generation
Intention Query Generation Input
You are an expert in generating queries for a target movie. Please help me generate the most suitable query for the target movie within one sentence, following the given example.
Example: TARGET: BUG-A-SALT 3.0 Black Fly Edition. QUERY: I want a gun that I can use while gardening to get rid of stink bugs, ants, flies, and spiders in my house. It needs to be amazing and help me feel less scared. TARGET: Toy Story (1995). Output
QUERY: I’m looking for a heartwarming animated movie that follows the adventures of a group of toys who come to life when their owner is not around.
Figure 7: Example of item query generation.
The user intention query is a natural language sentence implying the target item of interest. For each item in the dataset, we generate a fixed user intention query. Following the previous work [40], we generate user intention queries with the help of ChatGPT [88]. As shown in Figure 7, we prompt ChatGPT in a Chain-of-Thought (CoT) [103] paradigm and adopt the output as the user intention query. We adopt a rule-based strategy to ensure the quality of generated queries, and regenerate the wrong query. Considering the huge amount of item title text, we use ChatGPT3.5 API for generating all queries for the budget’s sake.
C.4.2 Baseline
AlphaRec exhibits user intention capture abilities, although not specially designed for search tasks. We compare AlphaRec with TEM [89] which falls in the field of personalized search [87, 104].
  • TEM [89] uses a transformer to encode the intention query together with user history behaviors, which enables it to achieve better search results by considering the user’s historical interest.
C.4.3 Case Study
We conduct two more case studies to verify the user intention capture ability of AlphaRec. As illustrated in Figure 8 and Figure 9, AlphaRec provides better recommendation results, assigning the target item at the top while maintaining the general user preferences.
Figure 8: Case study of user intention capture on MovieLens-1M
Figure 9: Case study of user intention capture on Video Games
C.4.4 Effect of the Intention Strength Alpha
The value of $\alpha$ controls the balance between the user’s historical interests and the user intention query. A larger $\alpha$ incorporates more about the user intention while considering less about the user’s historical interests. As shown in Figure 10, the effect of $\alpha$ on Video Games shows a similar trend with MovieLens-1M.
Figure 10: Effect of $\alpha$ on Video Games
C.5 Trainig Cost
We report the training cost of AlphaRec in this section. Table 10 reports the seconds needed per epoch and the total training cost until convergence. Here Amazon-Mix denotes the mixed dataset of Books, Movies & TV, and Video Games. It’s worth noting that AlphaRec converges quickly and only requires a small amount of training time.
Table 10: Training cost of AlphaRec (seconds per epoch/in total).
Books Movies & TV Video Games Amazon-Mix
AlphaRec 40.1 / 1363.4 12.3 / 479.7 7.4 / 214.6 107.2 / 5788.8
Appendix D Hyperparameter Settings and Implementation Details
Table 11: Hyperparameters search space for baselines.
Hyperparameter space
[…]
BC Loss $\tau_{1}$ $\sim$ [0.05, 3], $\tau_{2}$ $\sim$ [0.05, 2]
[…]
RLMRec kd weight $\sim$ [0.05, 2], kd temperature $\sim$ [0.01, 0.05, 0.1, 0.15, 0.2, 0.5, 1]
[…]
AlphaRec $\tau$ $\sim$ [0.05, 2]
Table 12: The hyperparameters of AlphaRec
Books Movies & TV Video Games Amazon-Mix
$\tau$ 0.15 0.15 0.2 0.15
We conduct all the experiments in PyTorch with a single NVIDIA RTX A5000 (24G) GPU and a 64 AMD EPYC 7543 32-Core Processor CPU. We optimize all methods with the Adam optimizer. For all ID-based CF methods, we set the layer numbers of graph propagation by default at 2, with the embedding size as 64 and the size of sampled negative items $|\mathcal{S}_{u}|$ as 256. We use the early stop strategy to avoid overfitting. We stop the training process if the Recall@20 metric on the validation set does not increase for 20 successive evaluations. In AlphaRec, the dimensions of the input and output in the two-layer MLP are 3072 and 64 respectively, with the hidden layer dimension as 1536. We apply the all-ranking strategy [105] for all experiments, which ranks all items except positive ones in the training set for each user. We search hyperparameters for baselines according to the suggestion in the literature. The hyperparameter search space is reported in Table 11. For these LM-enhanced models, KAR and RLMRec, we also search the hyperparameter of their backbone XSimGCL.
For AlphaRec, the only hyperparameter is the temperature $\tau$ and we search it in [0.05, 2]. We report the temperature $\tau$ we used for each dataset in Table 12. For the mixed dataset Amazon-Mix in Section 4.2, we use a universal $\tau$ = 0.15. We adopt $\tau$ = 0.2 for the MovieLens-1M dataset for the user intention capture experiment in Section 4.3.
Appendix E Broader Impact
The proposed AlphaRec can significantly improve the performance of zero-shot recommendation and the capability of user intent capture, offering a good approach to crafting more personalized recommendation results. One concern of AlphaRec is the potential for the representations generated by language models can be maliciously attacked, which may result in erroneous or unexpected recommendations. Therefore, we kindly advise researchers to cautiously check the quality of the language representations before using AlphaRec.

### v2

@@
-Language Models Encode Collaborative Signals in Recommendation
+Do Language Models Encode Collaborative Signals in Recommendation? Findings and Potentials
 Abstract
@@
 However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information.
-Contrary to the prevailing understanding that LMs and traditional recommender models learn two distinct representation spaces due to a huge gap in language and behavior modeling objectives, this work rethinks such understanding and explores extracting a recommendation space directly from the language representation space.
+Contrary to prevailing understanding that LMs and traditional recommenders learn two distinct representation spaces due to the huge gap in language and behavior modeling objectives, this work re-examines such understanding and explores extracting a recommendation space directly from the language representation space.
 Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance.
-This outcome suggests the homomorphism between the language representation space and an effective recommendation space, implying that collaborative signals may indeed be encoded within advanced LMs.
-Motivated by these findings, we propose a simple yet effective collaborative filtering (CF) model named AlphaRec, which utilizes language representations of item textual metadata (e.g., titles) instead of traditional ID-based embeddings.
-Specifically, AlphaRec is comprised of three main components: a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function, making it extremely easy to implement and train.
-Our empirical results show that AlphaRec outperforms leading ID-based CF models on multiple datasets, marking the first instance of such a recommender with text embeddings achieving this level of performance.
-Moreover, AlphaRec introduces a new language-representation-based CF paradigm with several desirable advantages: being easy to implement, lightweight, rapid convergence, superior zero-shot recommendation abilities in new domains, and being aware of user intention.
-Codes are available at https://github.com/LehengTHU/AlphaRec.
+This outcome suggests the possible homomorphism between the advanced language representation space and an effective item representation space for recommendation, implying that collaborative signals may be implicitly encoded within LMs.
+Motivated by the finding of homomorphism, we explore the possibility of designing advanced collaborative filtering (CF) models purely based on language representations without ID-based embeddings.
+To be specific, we incorporate several crucial components (i.e., a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function) to build a simple yet effective model, with the language representations of item textual metadata (i.e., title) as the input.
+Empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets, which sheds light on using language representations for better recommendation.
+Moreover, we systematically analyze this simple model and find several key features for using advanced language representations: a good initialization for item representations, superior zero-shot recommendation abilities in new datasets, and being aware of user intention.
+Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities  [^1] Codes are available at https://github.com/LehengTHU/AlphaRec.
 1 Introduction
-(a) Linearly mapping language representations into the recommendation space
+(a) Linearly mapping language representations into the behavior space for recommendation
 (b) Performance comparison
 (c) The t-SNE representations of movies and user intention in two spaces.
-Figure 1: Linearly mapping item titles in language representation space into recommendation space yields superior recommendation performance on Movies & TV [1] dataset.
+Figure 1: Linearly mapping item titles in language representation space into behavior space yields superior recommendation performance on Movies & TV (Ni et al., 2019) dataset.
 (1(a)) The framework of linear mapping.
 (1(b)) The recommendation performance comparison between leading CF recommenders and linear mapping.
-(1(c)) The t-SNE [2] visualizations of movie representations, with colored lines linking identical movies or user intention across language space (left) and linearly projected recommendation space (right).
-Language models (LMs) have achieved great success across various domains [3, 4, 5, 6, 7], prompting a critical question about the knowledge encoded within their representation spaces.
-Recent studies empirically find that LMs extend beyond semantic understanding to encode comprehensive world knowledge about various domains, including game states [8], lexical attributes [9], and even concepts of space and time [10] through language modeling.
-However, in the domain of recommendation where the integration of LMs is attracting widespread interest [11, 12, 13, 14, 15], it remains unclear whether LMs inherently encode relevant information on user preferences and behaviors.
-One possible reason is the significant difference between the objectives of language modeling for LMs and user behavior modeling for recommenders [16, 17, 18, 19].
-Currently, one prevailing understanding holds that general LMs and traditional recommenders encode two distinct representation spaces: the language space and the recommendation space (i.e., user and item representation space), each offering potential enhancements to the other for recommendation tasks [17, 20].
-On the one hand, when using LMs as recommenders, aligning the language space with the recommendation space could significantly improve the performance of LM-based recommendation [14, 21, 22, 23].
-Various alignment strategies are proposed, including fine-tuning LMs with recommendation data [15, 16, 24, 25, 26], incorporating embeddings from traditional recommenders as a new modality of LMs [17, 20, 27], and extending the vocabulary of LMs with item tokens [18, 19, 28, 29, 30, 31].
-On the other hand, when using LMs as the enhancer, traditional recommenders greatly benefit from from leveraging text representations [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], semantic and reasoning information [46, 47, 48, 49], and generated user behaviors [50, 51].
-Despite these efforts, explicit explorations of the relationship between language and recommendation spaces remain largely unexplored.
-In this work, we rethink the prevailing understanding and explore whether LMs inherently encode user preferences through language modeling.
-Specifically, we test the possibility of directly deriving a recommendation space from the language representation space, assessing whether the representations of item textual metadata (e.g., titles) obtained from LMs can independently achieve satisfactory recommendation performance.
-Positive results would imply that user behavioral patterns, such as collaborative signals (i.e., user preference similarities between items) [52, 53], may be implicitly encoded by LMs.
-To test this hypothesis, we employ linear mapping to project the language representations of item titles into a recommendation space (see Figure 1(a)).
-Our observations include:
-  • Surprisingly, this simple linear mapping yields high-quality item representations, which achieve exceptional recommendation performance (see Figure 1(b) and experimental results in Section 2).
-  • The clustering of items is generally preserved from the language space to the recommendation space (see Figure 1(c)).
-For example, movies with the theme of superheroes and monsters are gathering in both language and recommendation spaces.
-  • Interestingly, the linear mapping effectively reveals preference similarities that may be implicit or even obscure in the language space.
-For instance, while certain movies, such as those of homosexual movies (illustrated in Figure 1(c)), show dispersed representations in the language space, their projections through linear mapping tend to cluster together, reflecting their genres affiliation.
-These findings indicate the homomorphism [54] between the language representation space of LMs and an effective item representation space for recommendation.
-Motivated by this insight, we propose a new language-representation-based recommendation paradigm for general collaborative filtering (CF), which utilizes the pre-trained language representations of item titles as the item input and the average historical interactions’ representations as the user input.
-Different from traditional ID-based CF models [55, 56, 52] that heavily rely on trainable user and item IDs, this paradigm solely uses pre-trained LM embeddings and completely abandons ID-based embeddings.
-In this paper, to fully explore the potential of advanced language representations, we adopt a simple model architecture consisting of a two-layer MLP with graph convolution, and the popular contrastive loss, InfoNCE [57, 58, 59], as the objective function.
-This model is named AlphaRec for its originality, a series of good properties, and an adjustable hyperparameter $\alpha$ for user intention capture.
-Benefiting from paradigm shifts from ID-based embeddings to language representations, AlphaRec presents three desirable advantages.
-First, AlphaRec is notable for its simplicity, lightweight, rapid convergence, and exceptional recommendation performance (see Section 4.1).
-We empirically demonstrate that, for the first time, such a simple model with embeddings from pre-trained LMs can outperform leading CF models on multiple datasets.
-This finding strongly supports the possibility for developing language-representation-based recommender systems.
-Second, AlphaRec exhibits a strong zero-shot recommendation capability across untrained datasets (see Section 4.2).
-By co-training on three Amazon datasets (Books, Movies & TV, and Video Games) [1], AlphaRec can achieve performance comparable to the fully-trained LightGCN on entirely different platforms (MovieLens-1M [60] and BookCrossing [61]), and even exceed LightGCN in a completely new domain (Amazon Industrial), without additional training on these target datasets.
-This capability underscores AlphaRec’s potential to develop more general recommenders.
-Third, AlphaRec is user-friendly, offering a new research paradigm that enhances recommendation by leveraging language-based user feedback (see Section 4.3).
-Endowed with its inherent semantic comprehension of language representations, AlphaRec can refine recommendations based on user intentions expressed in natural language, enabling traditional CF recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
+(1(c)) The t-SNE (Van der Maaten & Hinton, 2008) visualizations of movie representations, with colored lines linking identical movies or user intention across language space (left) and linearly projected behavior space for recommendation (right).
+Language models (LMs) have achieved great success across various domains (Vaswani et al., 2017; Devlin et al., 2019; Dubey et al., 2024; OpenAI, 2023), raising a critical question about the knowledge encoded within the language space.
+Recent studies empirically find that LMs extend beyond semantic understanding to encode comprehensive world knowledge about various domains, such as game states (Li et al., 2023a), lexical attributes (Vulic et al., 2020), and even concepts of space and time (Gurnee & Tegmark, 2023) through language modeling.
+However, in the domain of recommendation where the integration of LMs is attracting widespread interest (Fan et al., 2023; Li et al., 2023b; Wu et al., 2023a), it remains unclear whether LMs inherently encode relevant information on user preferences and behaviors in the language space.
+Currently, one prevailing understanding holds that general LMs and traditional recommenders (e.g., collaborative filtering models (Koren et al., 2009; He et al., 2021)) encode distinct representation spaces — one for language space and the other for behavior space — but they offer the potential to enhance each other in downstream recommendation tasks (Liao et al., 2024).
+Specifically, on the one hand, when using LMs as recommenders to directly output items of interest, aligning the language space with the behavior space can significantly improve the recommendation performance (Lin et al., 2023a; Vats et al., 2024; Xu et al., 2024).
+Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023d; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024).
+On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024a), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; c).
+Despite these efforts, explicit studies of the relationship between language and behavior spaces remain largely unexplored in the recommendation domain.
+In this work, we re-examine this prevailing understanding, by exploring whether LM-generated language space has inherently encoded user preferences and behaviors.
+Specifically, we test the possibility of directly deriving a behavior space from the language space — that is, we assess whether the language representations of item text metadata (e.g., titles) generated by LMs can independently predict user behaviors and achieve competitive recommendation performance.
+Positive results would imply that user behavioral patterns, such as collaborative signals (i.e., users’ preference on items reflected by their behavioral similarities) (Wang et al., 2019b), might be implicitly encoded by LMs.
+To test this hypothesis, we employ linear mapping (Merullo et al., 2023) to project language representations of item titles into a behavior space for recommendation, as Figure 1(a) shows.
+Our empirical observations and findings include:
+  • Before linear mapping, language representation similarities (i.e., semantic textual similarities (STS) (Muennighoff et al., 2023)) may reflect user preference similarities for item contents.
+Considering Figure 1(c) as an example, movies with themes of superheroes and monsters cluster together in both language and behavior spaces.
+  • After linear mapping, language representations are transformed into high-quality behavior representations, which achieve exceptional recommendation performance, as Figure 1(b) and experimental results in Section 3.2 show.
+Moreover, the performance improves as the language model size increases and remains relatively robust to prompt disturbances (see Section 3.2).
+  • Post-mapping language representations encode user behaviroal similarities beyond STS.
+For instance, while certain movies, such as those of homosexual movies (illustrated in Figure 1(c)), show low STS and their representations disperse in the language space, their projections through linear mapping tend to cluster together, reflecting high user preference similarities.
+These findings suggest the homomorphism (Dieudonne, 1969) between the LM-generated language space and an expressive behavior space for recommendation.
+Motivated by this insight, we explore the possibility of building advanced collaborative filtering (CF) models based solely on the language space.
+To be specific, considering text metadata solely as items’ pre-existing features rather than the widely-used ID information, we perform a frozen LM to create the language representations of items; consequently, we apply a trainable projector (i.e., a two-layer MLP with graph convolution) to map them into a behavior space, and then employ a contrastive loss (i.e., InfoNCE (van den Oord et al., 2018; Wu et al., 2022)) to optimize.
+We term this model AlphaRec for its simplicity and a series of good properties.
+Surprisingly, our empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets.
+This result sheds light on using language representations for better recommendation.
+Furthermore, we systematically analyze this simple model and discover several potentials of adopting language representations for recommendation.
+First, language representations may serve as a good initialization for item representations, with few adjustments to achieve high recommendation performance (see Section 5.1).
+This is evidenced by the rapid training convergence of AlphaRec.
+Second, advanced language representations provide strong zero-shot recommendation capability across entirely new datasets (see Section 5.2).
+By co-training on multiple datasets, AlphaRec can achieve performance comparable to or even surpassing the fully-trained LightGCN (He et al., 2021) on new datasets without additional training.
+This underscores the potential of adopting advanced language representations to develop more generalizable recommenders.
+Third, advanced language representations provide opportunities for perceiving user intentions to refine recommendation results (see Section 5.3).
+Endowed with the inherent semantic comprehension of language representations, AlphaRec can adjust recommendations according to text-based user intentions, enabling recommenders to evolve into intention-aware systems through a straightforward paradigm shift.
+2 Preliminary
+2.1 Task Formulation
+Personalized recommendation aims to learn user preferences from historical behaviors (i.e., historical interactions with items like view, click, purchase) and find items of interest to trigger users’ future behaviors (Zhou et al., 2018).
+In this paper, we consider one common recommendation setting: collaborative filtering (CF) (Koren et al., 2022).
+It aims to select item $i\in\mathcal{I}$ that best matches user $u$’s preferences based on binary interaction behaviors $\mathbf{Y}=[y_{ui}]$, where $y_{ui}=1$ indicates user $u\in\mathcal{U}$ has interacted with item $i$, and $y_{ui}=0$ otherwise (Rendle, 2022).
+Scrutinizing leading CF models, we summarize a common paradigm $\hat{y}_{ui}=s\circ\phi_{\theta}(\mathbf{x}_{u},\mathbf{x}_{i})$ involving three components:
+  • For a user-item pair $(u,i)$, we first get their pre-existing features $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$, which are usually set as ID information or one-hot encodings in CF (Koren et al., 2009; Rendle, 2022; He et al., 2021).
+  • Upon $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$, the representation generation module $\phi$ parameterized by $\theta$ is adopted to transfer them into behavior representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, encoding the behavioral patterns of users.
+Its architecture can vary widely, including ID-based embeddings (Koren et al., 2009), multilayer perceptions (He et al., 2017), graph neural networks (Wang et al., 2019b; Cai et al., 2023), and variational autoencoders (Liang et al., 2018).
+  • Upon $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$, the scoring function $s$ is used to quantify their relevance reflecting how likely user $u$ will interact with item $i$.
+One widely-used function is cosine similarity, $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$ (Chen et al., 2023; Wu et al., 2022).
+2.2 Item Representation Generation
+Here we emphasize the critical role of item representation generation, which involves transforming item $i$’s pre-existing features $\mathbf{x}_{i}$ into representations $\mathbf{e}_{i}$ suitable for recommendation.
+This process is essential, as the quality of these representations directly impacts the recommendation performance.
+In this paper, we focus mainly on two kinds of item representation generation tailor-made for different pre-existing features: ID- and LM-based generators.
+ID-based generator.
+Prevailing CF models (Koren et al., 2009; Rendle, 2022; Koren et al., 2022; Wang et al., 2019b; He et al., 2021; Yu et al., 2024) typically convert the ID information of each item $i$ into one-hot encodings (e.g., pre-existing features $\mathbf{x}_{i}$).
+These sparse features are then passed through a trainable generator, such as ID embedding matrices (Koren et al., 2009; Rendle, 2022) or optionally combined with graph convolution layers (Wang et al., 2019b; He et al., 2021; Yu et al., 2024), to generate dense representations $\mathbf{e}_{i}$.
+Optimizing the learning of ID-based representations allows the generator to effectively learn user preferences and behaviors, leading to competitive recommendation performance.
+However, such ID-based generators suffer several problems, such as poor domain transferability and lack of user intention-aware abilities, since one-hot encodings lack sufficient semantics beyond being identifiers (He et al., 2021).
+LM-based generator.
+Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024b; Zhang et al., 2024a; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$.
+These features are fed into the LM-based generator, typically a combination of two subsequent components: (1) A frozen LM to extract $i$’s language representation $\mathbf{z}_{i}$, such as the encoder-only LMs like BERT-style models (Devlin et al., 2019; Liu et al., 2019), the decoder-only LMs like Llama-style autoregressive models (Touvron et al., 2023b; Jiang et al., 2023) and OpenAI text embedding models (Neelakantan et al., 2022); and (2) A trainable projector to map $\mathbf{z}_{i}$ into the final representation $\mathbf{e}_{i}$, often using layers like graph convolution layers (He et al., 2021).
+Although such LM-based generators have been explored to enrich the item representations in literatures (Yuan et al., 2023; Li et al., 2023c; Ren et al., 2024b), few studies have demonstrated that they can solely outperform ID-based generators in recommendation tasks.
+Worse still, the relationship between the LM-based language space and the behavior space remain largely unexplored in the recommendation domain.
 3 Uncovering Collaborative Signals in LMs via Linear Mapping
-In this section, we aim to explore whether LMs implicitly encode user preference similarities in their representation spaces.
-We first formulate the personalized item recommendation task, then detail the linear mapping method and its empirical findings.
-Empirical evidence indicates the homomorphism [54] between the representation spaces of advanced LMs and effective recommendation spaces.
-Task formulation.
-Personalized item recommendation with implicit feedback aims to select items $i\in\mathcal{I}$ that best match user $u$’s preferences based on binary interaction data $\mathbf{Y}=[y_{ui}]$, where $y_{ui}=1$ ($y_{ui}=0$) indicates user $u\in\mathcal{U}$ has (has not) interacted with item $i$ [59].
-The primary objective of recommendation is to model the user-item interaction matrix $\mathbf{Y}$ using a scoring function $\hat{y}:\mathcal{U}\times\mathcal{I}\rightarrow\mathbb{R}$, where $\hat{y}_{ui}$ measures $u$’s preference for $i$.
-The scoring function $\hat{y}_{ui}=s\circ\phi_{\theta}(\mathbf{x}_{u},\mathbf{x}_{i})$ comprises three key components: pre-existing features $\mathbf{x}_{u}$ and $\mathbf{x}_{i}$ for user $u$ and item $i$, a representation learning module $\phi_{\theta}(\cdot,\cdot)$ parametrized by $\theta$, and a similarity function $s(\cdot,\cdot)$.
-The representation learning module $\phi_{\theta}$ transfers $u$ and $i$ into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ for similarity matching $s(\cdot,\cdot)$, and the Top-$K$ highest scoring items are recommended to $u$.
-Different recommenders employ various pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$ and representation learning architecture $\phi_{\theta}(\cdot,\cdot)$.
-Traditional ID-based recommenders use one-hot vectors as pre-existing features $\mathbf{x}_{u},\mathbf{x}_{i}$.
-The choice of ID-based representation learning architecture $\phi_{\theta}$ can vary widely, including ID-based embedding matrix [55], multilayer perception [62], graph neural network [52, 63], and variational autoencoder [64].
-We adopt a commonly used similarity function, cosine similarity [65, 58] $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, in this paper.
-Linear mapping.
-Linear mapping ensures the homomorphism [54] between the vector spaces, and is therefore widely used in the study on the representation space properties of LMs [10, 66, 67].
-In this section, we adopt a trainable linear mapping matrix $\bm{W}$ as the representation learning module $\phi_{\theta}$, mapping representations from the language space into an item recommendation space.
-High performance on the test set would indict the homomorphism between the language space and an effective recommendation space, which further suggests the possible existence of collaborative signals (i.e., user preference similarities between items) [56] in the language representation space [68, 10].
-The overall framework of the linear mapping process is illustrated in Figure 1(a).
-Specifically, we utilize LMs as feature extractors, leveraging the language representations of item titles as initial item feature $\mathbf{x}_{i}$.
-For initial user feature $\mathbf{x}_{u}$, we use the average of the title representations of historically interacted items, defined as $\mathbf{x}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{x}_{i}$, where $\mathcal{N}_{u}$ is the set of items user $u$ has interacted with.
-Detailed procedures for obtaining these language-based features are provided in Appendix B.2.
-The linear mapping matrix sets user and item representations as $\mathbf{e}_{u}=\bm{W}\mathbf{x}_{u}$ and $\mathbf{e}_{i}=\bm{W}\mathbf{x}_{i}$ respectively.
-To learn the linear mapping matrix $\bm{W}$, we adopt the InfoNCE loss [57] as the objective function, which has demonstrated state-of-the-art performance in both ID-based [69, 70] and LM-enhanced CF recommenders [47] (refer to Equation (4) for the formula).
-We directly use linearly mapped representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ to calculate the user-item similarity $s(\mathbf{e}_{u},\mathbf{e}_{i})$ for recommendation.
-Empirical findings.
-We compare the recommendation performance of the linear mapping method with three classical CF baselines, matrix factorization (MF) [55, 71], MultVAE [64], and LightGCN [56] (see more details about baselines in Appendix C.2.1).
-We report three widely used metrics Hit Ratio (HR@$K$), Recall@$K$, Normalized Discounted Cumulative Gain (NDCG@$K$)) to evaluate the effectiveness of linear mapping, with $K$ set by default at 20.
-We evaluate a wide range of LMs, including BERT-style models [4, 5], decoder-only large language models (LLMs) [6, 72], and LLM-based text embedding models [73, 74] (see Appendix B.1 for details about used LMs).
-Table 1: The recommendation performance of linear mapping comparing with classical CF baselines.
-Books Movies & TV Video Games
+In this section, we first explore the following research questions.
+RQ1: Do LMs inherently encode collaborative signals (i.e., users’ preferences for items as reflected by behavioral similarities) within their representation spaces?
+RQ2: If so, does the presence of such signals scale with model size, and are they robust across different settings?
+To investigate these questions, we use linear mapping to project language representations of item titles into a behavior space for recommendation.
+We detail the implementation of the linear mapping in Section 3.1.
+Subsequently, in Section 3.2, we empirically assess the existence and robustness of collaborative signals in language representations.
+3.1 Linear Mapping
+Linear mapping is effective to study the representation properties of LMs (Merullo et al., 2023; Alain & Bengio, 2017), discovering the homomorphism (Dieudonne, 1969) between the language space and another space in the target domain.
+However, its application in the recommendation domain remains largely underexplored.
+To bridge this gap, we train a linear mapping matrix $\bm{W}$ to project representations from the language space into a behavior space for recommendation.
+High performance of this linear mapping on the test set would indicate the presence of homomorphism between the language space and an effective behavior space, suggesting the possible existence of collaborative signals in the language representation space (Ravichander et al., 2021; Gurnee & Tegmark, 2023).
+The overall framework of linear mapping is illustrated in Figure 1(a).
+Specifically, we use frozen LMs to transform pre-existing item title features $\mathbf{x}_{i}$ into language representations $\mathbf{z}_{i}$.
+To derive user representations, we compute the average of the language representations of the items a user $u$ has interacted with, denoted as $\mathbf{z}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{z}_{i}$, where $\mathcal{N}_{u}$ is the set of user $u$’s historical items.
+See Appendix B.2 for detailed procedures for obtaining language representations.
+The linear mapping matrix sets behavior representations of user $u$ and item $i$ as $\mathbf{e}_{u}=\bm{W}\mathbf{z}_{u}$ and $\mathbf{e}_{i}=\bm{W}\mathbf{z}_{i}$ respectively.
+To optimize the matrix $\bm{W}$, we adopt the InfoNCE loss (van den Oord et al., 2018) as the objective function, which has demonstrated strong performance in both ID-based (Zhang et al., 2023a; Yu et al., 2024) and LM-based generators (Ren et al., 2024b) (refer to equation 4 for the formula).
+3.2 Empirical Findings
+Table 1: The comparison of the recommendation performance of linear mapping with the classical ID-based CF baselines.
+Movies & TV Video Games Books
 Recall NDCG HR Recall NDCG HR Recall NDCG HR
-CF MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
-MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $\underline{0.3418}$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
-LightGCN (He et al., 2021) $0.0723$ 0.0608 0.3489 $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
-Linear Mapping BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
-RoBERTa 0.0247 0.0209 0.1262 0.0406 0.0387 0.2277 0.0578 0.0338 0.1339
-Llama2-7B 0.0662 0.0559 0.3176 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746
-Mistral-7B 0.0650 0.0544 0.3124 0.1039 0.0963 0.4994 0.1270 0.0687 0.2428
-text-embedding-ada-v2 0.0515 0.0436 0.2570 0.0926 0.0874 0.4563 0.1176 0.0683 0.2579
-text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
-SFR-Embedding-Mistral 0.0738 0.0610 0.3371 0.1152 0.1065 0.5327 0.1370 0.0787 0.2927
-Table 1 reports the recommendation performance yielded by the linear mapping on three Amazon datasets [1], comparing with classic CF baselines.
-We observe that the performance of most advanced text embedding models (e.g., text-embeddings-3-large [73] and SFR-Embedding-Mistral [74]) exceeds LightGCN on all datasets.
-We also empirically show that these improvements may not merely come from the better feature encoding ability (refer to Appendix B.3).
-Figure 1(c) further suggests that obscure user preference similarities may be implicitly encoded in the language representation space, with a linear mapping matrix to activate.
-These findings indicate the homomorphism [54] between the language representation spaces of advanced LMs and effective item representation spaces for recommendation.
-In addition, with the help of linear mapping, for the first time, we are able to clearly quantify the gap in the representation spaces between different LMs for recommendation.
-With the advances in LMs, we observe that the performance of language representations exhibits a rising trend, especially showing a sharp rise from BERT-style models to LLMs.
-Representations from early BERT-style models (e.g., BERT [4] and RoBERTa [5]) only show weaker or equal capabilities compared with MF, failing behind LLMs largely.
-This further suggests that, until very recently, the language representation spaces are able to encode user preference similarities efficiently.
-3 AlphaRec
-This finding of space homomorphism [54] sheds light on building advanced CF models purely based on language representations without introducing ID-based embeddings.
-Inspired by this, we develop a simple yet effective CF model called AlphaRec, by incorporating only three simple components (i.e., nonlinear projection [62], graph convolution [56] and contrastive learning (CL) objectives [57]).
-It is important to highlight that our approach is centered on exploring the potential of language representations for recommendation by integrating essential components from leading CF models, rather than deliberately inventing new CF mechanisms.
-Surprisingly, these basic and simple components lead to an excellent language-representation-based recommender with multiple capabilities.
-We present the model structure of AlphaRec in Section 3.1, and compare AlphaRec with two popular recommendation paradigms in Section 3.2.
-3.1 Method
-We present how AlphaRec is designed and trained.
-Generally, the representation learning architecture $\phi_{\theta}(\cdot,\cdot)$ of AlphaRec is simple, which only contains a two-layer MLP and the basic graph convolution operation, with language representations as the input features $\mathbf{x}_{u},\mathbf{x}_{i}$.
-The cosine similarity is used as the similarity function $s(\cdot,\cdot)$, and the contrastive loss InfoNCE [57, 58] is adopted for optimization.
-For simplicity, we consistently adopt text-embeddings-3-large [73] as the language representation model, for its excellent language understanding and representation capabilities.
+CF MF (Rendle et al., 2012) $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$ $0.0437$ $0.0391$ $0.2476$
+MultVAE (Liang et al., 2018) $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$ $0.0722$ $0.0597$ $\underline{0.3418}$
+LightGCN (He et al., 2021) $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$ $0.0723$ 0.0608 0.3489
+Linear Mapping BERT 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245 0.0226 0.0194 0.1240
+RoBERTa 0.0406 0.0387 0.2277 0.0578 0.0338 0.1339 0.0247 0.0209 0.1262
+Llama2-7B 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746 0.0662 0.0559 0.3176
+Mistral-7B 0.1039 0.0963 0.4994 0.1270 0.0687 0.2428 0.0650 0.0544 0.3124
+text-embedding-ada-v2 0.0926 0.0874 0.4563 0.1176 0.0683 0.2579 0.0515 0.0436 0.2570
+text-embeddings-3-large 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928 0.0735 0.0608 0.3355
+SFR-Embedding-Mistral 0.1152 0.1065 0.5327 0.1370 0.0787 0.2927 0.0738 0.0610 0.3371
+(a) Movie & TV
+(b) Games
+(c) Books
+Figure 2: The recommendation performance of linear mapping with different language model sizes.
+Existence (RQ1).
+To explore the existence of collaborative signals in language representations, we test the recommendation performance of the linear mapping method.
+Table 1 reports the performance yielded by post-mapping representations on three Amazon datasets (Ni et al., 2019), comparing with classic ID-based CF baselines: matrix factorization (MF) (Koren et al., 2009), MultVAE (Liang et al., 2018), and LightGCN (He et al., 2021) (see baseline details in Appendix C.2).
+Figures (2(a)) - (2(c)) depict the linear mapping performance under different LM sizes.
+Figure 1(c) demonstrates the visualization of representations before and after linear mapping.
+We observe that:
+  • Post-mapping representations of advanced LMs achieve superior recommendation performance in most cases, suggesting the possible homomorphism between language spaces and behavior spaces.
+Specifically, advanced LMs (e.g., Llama2-7B (Touvron et al., 2023b) and text-embeddings-3-large (Neelakantan et al., 2022)) consistently perform better than leading CF models (e.g., LightGCN) on most metrics.
+We also see that the performance improves with more recent and advanced LMs.
+In contrast, earlier BERT-style models (e.g., BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)) perform similarly to or worse than MF, indicating that LMs have only recently developed the ability to encode user preference similarities effectively.
+  • Language representations encode user preference similarities beyond semantic textual similarities (STS).
+Consider Figure 1(c) as an example again, homosexual movies, which differ significantly in textual meaning, cluster together after linear mapping.
+This suggests that user preferences, which are not immediately apparent from text alone, are implicitly encoded in the language space and can be uncovered through linear mapping.
+Table 2: The robustness of language representations for recommendation.
+Movies & TV Video Games Books
+Recall NDCG HR Recall NDCG HR Recall NDCG HR
+Title + Random Noise 0.0952 0.0887 0.4731 0.1213 0.0706 0.2722 0.0632 0.0525 0.3099
+Title Only 0.1027 0.0955 0.4952 0.1249 0.0729 0.2746 0.0662 0.0559 0.3176
+Scaling and Robustness (RQ2).
+We also investigate whether the encoding of user preference similarities in LMs scales with model size and whether this encoding is robust in the presence of noise in the input prompts.
+To this end, we test the linear mapping performance of LMs of various sizes and prompts with noise:
+  • The encoding of user preference similarities becomes more refined as model size increases, leading to better linear mapping performance.
+Specifically, we test the linear mapping performance across different language model sizes (7B, 13B, and 70B) of the Llama2 family (Touvron et al., 2023b).
+Llama3 (Dubey et al., 2024) is not selected due to the lack of a 13B model.
+As shown in Figure 2, linear mapping performance improves consistently as the model size increases from 7B to 70B, indicating that larger models capture more nuanced user behavioral patterns.
+  • Language representations are relatively robust to prompt disturbances.
+Following previous works (Gurnee & Tegmark, 2023), we compare two prompting strategies: using item titles alone (e.g., Castlevania), and adding 5-10 random letters to the titles (e.g., Castlevania sdfhdsk).
+Table 2 shows that adding random noise to the item titles had minimal impact on the linear mapping performance.
+The prompt noise has more impact on Movies & TV since item titles in this dataset are relatively shorter than others (see Appendix C.1).
+This finding suggests the relative robustness of the recommendation knowledge encoded in the language representation space.
+4 Leveraging Language Representations for Better Recommendation
+This finding of possible space homomorphism (Dieudonne, 1969) and encoded collaborative signals arouse interest in the following questions.
+RQ3: How powerful are such language representations for building advanced CF models that can outperform prevailing ID-based CF methods?
+To address these questions, in Section 4.1, we aim to develop a simple yet effective CF model termed AlphaRec, which is solely based on language representations and merely incorporates three crucial components in modern CF models.
+After that, we evaluate its performance in Section 4.2 to demonstrate the capability of advanced language representations for recommendation.
+4.1 AlphaRec
+We briefly present how this simple model AlphaRec is designed and trained.
+It is important to highlight that we center on exploring the power of language representations for recommendation, rather than deliberately inventing new CF mechanisms.
+Generally, the representation generation architecture $\phi_{\theta}(\cdot,\cdot)$ is simple, which only contains a two-layer MLP and the basic graph convolution operation.
+The cosine similarity is used as the similarity function $s(\cdot,\cdot)$, and the contrastive loss InfoNCE (van den Oord et al., 2018; Wu et al., 2022) is adopted for optimization.
+For simplicity, we adopt text-embeddings-3-large (Neelakantan et al., 2022) for language representation generation by default, for its excellent language understanding and representation capabilities.
 Nonlinear projection.
-In AlphaRec, we substitute the linear mapping matrix delineated in Section 2 with a nonlinear MLP.
-This conversion from linear to nonlinear is non-trivial, for the paradigm shift from ID-based embeddings to language representations, since nonlinear transformation helps in excavating more comprehensive user preference similarities from the language representation space with rich semantics (see discussions about this in Appendix C.2.3) [62].
-Specifically, we project the language representation $\mathbf{x}_{i}$ of the item title to an item space for recommendation with the two-layer MLP, and obtain user representations as the average of historical items:
-$\mathbf{e}_{i}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{2},\quad\mathbf{e}_{u}^{(0)}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{e}^{(0)}_{i}.$ (1)
+We substitute the linear matrix delineated in Section 3 with a nonlinear MLP.
+Nonlinear transformation helps in excavating more comprehensive preference similarities from the language representation space (see discussions about this in Appendix C.4) (He et al., 2017).
+Taking the averaged language representaions of historical items as the user language representation (i.e., $\mathbf{z}_{u}=\frac{1}{|\mathcal{N}_{u}|}\sum_{i\in\mathcal{N}_{u}}\mathbf{z}_{i}$), the initial nonlinear transformation operation be formulated as:
+$\mathbf{e}_{i}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{z}_{i}+\bm{b}_{1}\right)+\bm{b}_{2},\quad\mathbf{e}_{u}^{(0)}=\bm{W}_{2}\operatorname{LeakyReLU}\left(\bm{W}_{1}\mathbf{z}_{u}+\bm{b}_{1}\right)+\bm{b}_{2}.$ (1)
 Graph convolution.
-Graph neural networks (GNNs) have shown superior effectiveness for recommendation [52, 56], owing to the natural user-item graph structure in recommender systems [75].
-In AlphaRec, we employ a minimal graph convolution operation [56] to capture more complicated collaborative signals from high-order connectivity [56, 76, 77, 75] as follows:
+Graph neural networks (GNNs) show superior effectiveness for recommendation (Wang et al., 2019b), owing to the natural user-item graph structure in recommender systems (Wu et al., 2023b).
+We employ a minimal graph convolution operation (He et al., 2021) to capture more complicated collaborative patterns from high-order connectivity (Wu et al., 2019) as follows:
 $\mathbf{e}_{u}^{(k+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{1}{\sqrt{\left|\mathcal{N}_{u}\right|}\sqrt{\left|\mathcal{N}_{i}\right|}}\mathbf{e}_{i}^{(k)},\quad\mathbf{e}_{i}^{(k+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{1}{\sqrt{\left|\mathcal{N}_{i}\right|}\sqrt{\left|\mathcal{N}_{u}\right|}}\mathbf{e}_{u}^{(k)}.$ (2)
@@
 Here $\mathcal{N}_{u}$ ($\mathcal{N}_{i}$) denotes the historical item (user) set that user $u$ (item $i$) has interacted with.
-The features $\mathbf{e}^{(0)}_{u}$ and $\mathbf{e}^{(0)}_{i}$ projected from the MLP are used as the input of the first layer.
-After propagating for $K$ layers, the final representation of a user (item) is obtained as the average of features from each layer:
+The representations $\mathbf{e}^{(0)}_{u}$ and $\mathbf{e}^{(0)}_{i}$ projected from the MLP are used as the input of the first layer.
+After propagating for $K$ layers, the final behavior representation of a user $u$ (item $i$) is obtained as the average of representations from each layer:
 $\mathbf{e}_{u}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{u}^{(k)},\quad\mathbf{e}_{i}=\frac{1}{K+1}\sum_{k=0}^{K}\mathbf{e}_{i}^{(k)}.$ (3)
 Contrastive learning objective.
-The introduction of contrasting learning is another key element for the success of leading CF models.
-Recent research suggests that the contrast learning objective, rather than data augmentation, plays a more significant role in improving recommendation performance [70, 78, 69].
-Therefore, we simply use the contrast learning object InfoNCE [57] as the loss function without any additional data augmentation on the graph [79, 58].
-With cosine similarity as the similarity function $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, the InfoNCE loss [57, 79, 80] is written as:
+The introduction of contrastive learning (Radford et al., 2021) is another key element for the success of leading CF models.
+Recent research suggests that the contrast learning objective, rather than data augmentation, plays a more significant role in improving recommendation performance (Yu et al., 2024; 2022; Zhang et al., 2023a).
+Therefore, we simply use the contrast learning object InfoNCE (van den Oord et al., 2018) as the loss function without any additional data augmentation on the graph (Wu et al., 2022).
+With cosine similarity as the similarity function $s(\mathbf{e}_{u},\mathbf{e}_{i})=\frac{{\mathbf{e}_{u}}^{\top}\mathbf{e}_{i}}{\left\lVert\mathbf{e}_{u}\right\rVert\cdot\left\lVert\mathbf{e}_{i}\right\rVert}$, the InfoNCE loss (van den Oord et al., 2018) is written as:
 $\mathbf{\mathop{\mathcal{L}}}_{\text{InfoNCE}}=-\sum_{(u,i)\in\mathcal{O}^{+}}\log\frac{\exp{(s(\mathbf{e}_{u},\mathbf{e}_{i})/\tau)}}{\exp{(s(\mathbf{e}_{u},\mathbf{e}_{i})/\tau)}+\sum_{j\in\mathcal{S}_{u}}\exp{(s(\mathbf{e}_{u},\mathbf{e}_{j})/\tau)}}.$ (4)
-Here, $\tau$ is a hyperparameter called temperature [81], $\mathcal{O}^{+}=\{(u,i)|y_{ui}=1\}$ denoting the observed interactions between users $\mathcal{U}$ and items $\mathcal{I}$.
+Here, $\tau$ is a hyperparameter called temperature (Wang & Liu, 2021), $\mathcal{O}^{+}=\{(u,i)|y_{ui}=1\}$ denoting the observed interactions between users $\mathcal{U}$ and items $\mathcal{I}$.
 And $\mathcal{S}_{u}$ is a randomly sampled subset of negative items that user $u$ does not adopt.
-3.2 Discussion of Recommendation Paradigms
-We compare the language-representation-based AlphaRec with two popular recommendation paradigms in Table 2 (see more discussion about related works in Appendix A).
-ID-based recommendation (ID-Rec) [52, 55].
-In the traditional ID-based recommendation paradigm, users and items are represented by ID-based learnable embeddings derived from a large number of user interactions.
-While ID-Rec exhibits excellent recommendation capabilities with low training and inference costs [63, 79], it also has two significant drawbacks.
-Firstly, these ID-based embeddings learned in specific domains are difficult to transfer to new domains without overlapping users and items [37], thereby hindering zero-shot recommendation capabilities.
-Additionally, there is a substantial gap between ID-Rec and natural languages [34], which makes ID-based recommenders hard to incorporate language-based user intentions and further refine recommendations accordingly.
-LM-based recommendation (LM-Rec) [15, 16, 24].
-Benefitting from the extensive world knowledge and powerful reasoning capabilities of LMs [7, 82], the LM-based recommendation paradigm has gained widespread attention [11, 13].
-LM-Rec tends to convert user interaction history into text prompts as input for LMs, utilizing pre-trained or fine-tuned LMs in a text generation pattern to recommend items.
-LM-Rec demonstrates zero-shot and few-shot abilities and can easily understand language-based user intentions.
-However, LM-Rec faces significant challenges.
-Firstly, the LM-based model architecture leads to huge training and inference costs, with real-world deployment difficulties.
-Additionally, limited by the text generation paradigm, LM-based models tend to perform candidate selection [17] or generate a single next item [24].
-It remains difficult for LM-Rec to comprehensively rank the entire item corpus or recommend multiple items that align with user interests.
-Language-representation-based recommendation.
-We argue that AlphaRec follows a new CF paradigm, which we term the language-representation-based paradigm.
-This paradigm replaces the ID-based embeddings in ID-Rec with representations from pre-trained LMs, employing feature encoders to map language representations directly into the recommendation space.
-Few early studies lie in this paradigm, including using BERT-style LMs to learn universal sequence representations [37, 44], or adopting the same model architecture as ID-Rec with simple input features replacement [34, 35].
-These early explorations, which are mostly based on BERT-style LMs, are usually only applicable in certain specific scenarios, such as the transductive setting with the help of ID-based embeddings [37].
-This phenomenon is consistent with our previous findings in Section 2, indicating that BERT-style LMs may fail to effectively encode user preference similarities.
-We point out that AlphaRec is the first recommender in the language-representation-based paradigm to surpass the traditional ID-based paradigm on multiple tasks, faithfully demonstrating the effectiveness and potential of this paradigm.
-Table 2: Comparison of recommendation paradigms
-Recommendation Paradigms Training Cost Zero-shot Ability Intention-aware Ability
-ID-based Low
-LM-based High
-Language-representation-based Low
-4 Experiments
-In this section, we aim to explore the effectiveness of AlphaRec. Specifically, we are trying to answer the following research questions:
-  • RQ1: How does AlphaRec perform compared with leading ID-based CF methods?
-  • RQ2: Can AlphaRec learn general item representations, and achieve good zero-shot recommendation performance on entirely new datasets?
-  • RQ3: Can AlphaRec capture user intention described in natural language and adjust the recommendation results accordingly?
-4.1 General Recommendation Performance (RQ1)
-Motivation.
-We aim to explore whether the language-representation-based recommendation paradigm can outperform the ID-Rec paradigm.
-An excellent performance of AlphaRec would shed light on the research line of building language-representation-based recommenders in the future.
+4.2 Empirical Findings
 Baselines.
-We consider ID-based baselines in this section.
-LM-based methods are not incorporated for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection or next item prediction.
-In addition to classic baselines (i.e., MF, MultVAE, and LightGCN) introduced in section 2, we consider two categories of leading ID-based CF baselines: CL-based CF methods: SGL [83], BC Loss [79], XSimGCL [70] and LM-enhanced CF methods: KAR [48], RLMRec [47].
-See more details about baselines in Appendix C.2.1.
-Results.
-Table 3 presents the performance of AlphaRec compared with leading CF baselines.
-The best-performing methods are bold, while the second-best methods are underlined.
-Figure 2(a) and Figure 2(b) report the training efficiency and ablation results.
-We observe that:
+We compare AlphaRec with leading ID-based CF baselines, to assess the effectiveness of adopting advanced language representations.
+We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023).
+In addition to classic baselines introduced in section 3.2, we consider two categories of leading ID-based CF baselines, CL-based CF methods: SGL (Wu et al., 2021), BC Loss (Zhang et al., 2022), XSimGCL (Yu et al., 2024) and LM-enhanced CF methods: KAR (Xi et al., 2023), RLMRec (Ren et al., 2024b).
+See more details about baselines in Appendix C.2.
+Recommendation capabilities (RQ3). Table 3 presents performance comparison. The best-performing methods are bold, while the second-best methods are underlined. We observe that:
 Table 3: The performance comparison with ID-based CF baselines. The improvement achieved by AlphaRec is significant ($p$-value <<much-less-than$<$ < 0.05).
-Books Movies & TV Video Games
+Movies & TV Video Games Books
 Recall NDCG HR Recall NDCG HR Recall NDCG HR
-MF (Rendle et al., 2012) $0.0437$ $0.0391$ $0.2476$ $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$
-MultVAE (Liang et al., 2018) $0.0722$ $0.0597$ $0.3418$ $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$
-LightGCN (He et al., 2021) $0.0723$ $0.0608$ $0.3489$ $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$
-SGL (Wu et al., 2021) 0.0789 0.0657 0.3734 0.0916 0.0838 0.4680 0.1089 0.0634 0.2449
-BC Loss (Zhang et al., 2022) 0.0915 0.0779 0.4045 0.1039 0.0943 0.5037 0.1145 0.0668 0.2561
-XSimGCL (Yu et al., 2024) 0.0879 0.0745 0.3918 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550
-KAR (Xi et al., 2023) 0.0852 0.0734 0.3834 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571
-RLMRec (Ren et al., 2024a) 0.0928 0.0774 0.4092 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997
-AlphaRec 0.0991* 0.0828* 0.4185* 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207*
+MF (Rendle et al., 2012) $0.0568$ $0.0519$ $0.3377$ $0.0323$ $0.0195$ $0.0864$ $0.0437$ $0.0391$ $0.2476$
+MultVAE (Liang et al., 2018) $0.0853$ $0.0776$ $0.4434$ $0.0908$ $0.0531$ $0.2211$ $0.0722$ $0.0597$ $0.3418$
+LightGCN (He et al., 2021) $0.0849$ $0.0747$ $0.4397$ $0.1007$ $0.0590$ $0.2281$ $0.0723$ $0.0608$ $0.3489$
+SGL (Wu et al., 2021) 0.0916 0.0838 0.4680 0.1089 0.0634 0.2449 0.0789 0.0657 0.3734
+BC Loss (Zhang et al., 2022) 0.1039 0.0943 0.5037 0.1145 0.0668 0.2561 0.0915 0.0779 0.4045
+XSimGCL (Yu et al., 2024) 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550 0.0879 0.0745 0.3918
+KAR (Xi et al., 2023) 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571 0.0852 0.0734 0.3834
+RLMRec (Ren et al., 2024b) 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997 0.0928 0.0774 0.4092
+AlphaRec 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207* 0.0991* 0.0828* 0.4185*
 Imp.over the best baseline $6.79\%$ $5.34\%$ $2.27\%$ $9.12\%$ $10.75\%$ $5.40\%$ $9.75\%$ $10.51\%$ $7.01\%$
-(a) Training efficiency comparison
-(b) Ablation study on Books
-Figure 2: (2(a)) The bar charts show the number of epochs needed for each model to converge.
-AlphaRec tends to exhibit an extremely fast convergence speed.
-(2(b)) The effect of each component in AlphaRec on Books dataset.
-  • AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets.
-AlphaRec shows an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline RLMRec [47].
-We further conduct the ablation study to explore the reason for its success (see more ablation results in Appendix C.2.2).
-As shown in Figure 2(b), each component in AlphaRec contributes positively.
+(a) Ablation study on Books
+(b) Training efficiency comparison
+Figure 3: (3(a)) The effect of each component on Books dataset. (3(b)) The number of epochs needed for each model to converge. AlphaRec exhibits a breakneck convergence speed.
+  • Advanced language representations shows strong potentials for recommendation, which can be unleashed by appropriate model design.
+AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets, with an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline.
+Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results in Appendix C.3) in unleashing the power of language representations.
 Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively.
-Moreover, the performance drop from replacing InfoNCE loss [58] with BPR loss [71] (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance.
-These findings suggest that, by carefully designing the model to extract user preference similarities, the potential of language-representation-based paradigm can be realized, surpassing the leading ID-Rec paradigm.
-  • The incorporation of semantic language representations into traditional ID-based CF methods can lead to significant performance improvements.
-We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over CL-based CF methods.
+Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance.
+These findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
+  • The incorporation of advanced language representations can benefit traditional ID-based CF methods.
+We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over the most advanced CF methods.
 Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec.
-We attribute this phenomenon to the fact that the performance contribution of these methods mainly comes from the language representations, which is consistent with the previous findings [34, 44].
-  • AlphaRec exhibits fast convergence speed.
-We find that the convergence speed of AlphaRec is comparable with, or even surpasses, CL-based methods with data augmentation (e.g., SGL [83] and XSimGCL [70]).
-Meanwhile, methods based solely on graph convolution (LightGCN [56]) or CL objective (BC Loss [79]) show relatively slow convergence speed, indicating that introducing these modules may not lead to convergence speed improvement.
-We attribute the fast convergence speed of AlphaRec to the homomorphism between the language space and a good recommendation space.
-As a result, only minor adjustments to the vanilla language representations are needed for effective recommendation.
-4.2 Zero-shot Recommendation Performance on Entirely New Datasets (RQ2)
-Motivation. We aim to explore whether AlphaRec has learned general item representations [37], which enables it to perform well on entirely new datasets without any user and item overlap.
-Task and datasets.
-In zero-shot recommendation [38], there is not any item or user overlap between the training set and test set [38, 33], which is different from the research line of cross-domain recommendation in ID-Rec [84].
-We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M [60], Book Crossing [61], and Industrial [1]) without further training on these new datasets.
-(see more details about how we train AlphaRec on multiple datasets in Appendix C.3.1).
-Baselines.
-Due to the lack of zero-shot recommenders in the field of general recommendation, we slightly modify two zero-shot methods in the sequential recommendation [85], ZESRec [37] and UniSRec [37], as baselines.
-We also incorporate two strategy-based CF methods, Random and Pop (see more details about these baselines in Appendix C.3.2).
+We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024c).
+5 Exploring Potentials of Language Representations for Recommendation
+In this section, we focus on this question: What new opportunities beyond good performance can advanced language representations bring to recommender systems?
+To answer this, we systematically analyze AlphaRec and discover the following potentials of adopting such language representations.
+Potential 1: Good initialization for item representations (Section 5.1).
+Potential 2: Zero-shot ability (Section 5.2).
+Potential 3: Intention-aware ability (Section 5.3).
+5.1 Good Initialization for Item Representations (Potential 1)
+Advanced language representations may provide a good initialization for item representations, with few adjustments for effective recommendation.
+As shown in Figure 3(b), beyond its good performance, AlphaRec also exhibits extremely fast convergence speed, which is comparable with or even surpasses the fastest ID-based CF methods (e.g., SGL (Wu et al., 2021) and XSimGCL (Yu et al., 2024)).
+Moreover, recent works also suggest that, when using advanced language representations to initialize ID-based item embeddings (Harte et al., 2023; Zhao et al., 2024), the performance of traditional ID-based recommenders improves significantly.
+We attribute the above findings to the homomorphism between the language space and a good behavior space.
+Therefore, when using advanced language representations for initialization, only minor adjustments are needed to generate effective behavior representations for recommendation.
+5.2 Zero-shot Ability (Potential 2)
+The prevailing ID-based recommenders suffer from domain transferring problems (i.e., behavior representations are highly bound with ID information (Zhu et al., 2021)).
+Advanced language representations may provide opportunities for learning transferable item representations (Hou et al., 2022), enabling recommenders to perform well on entirely new datasets without any ID overlap.
+To address this potential, we test the zero-shot recommendation ability of AlphaRec (Ding et al., 2021).
+Experimental settings.
+In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024b), which is different from the research line of cross-domain recommendation (Zhu et al., 2021).
+We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M (Harper & Konstan, 2016), Book Crossing (Lee et al., 2019), and Amazon Industrial & Scientific (Ni et al., 2019)) without further training on these new datasets.
+(see more details about training on multiple datasets in Appendix D.2.1).
+Due to the lack of zero-shot recommenders in general CF, we slightly modify the two zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022) and UniSRec (Hou et al., 2022), as baselines.
+We also incorporate two strategy-based CF methods (i.e., Random and Pop) and one method using the large language model (LLM) as zero-shot recommender (i.e., LLMRank (Hou et al., 2024b)) (see more details about baselines in Appendix D.2.2).
 Table 4: The zero-shot recommendation performance comparison on entirely new datasets. The improvement achieved by AlphaRec is significant ($p$-value <<much-less-than$<$ < 0.05).
-Industrial MovieLens-1M Book Crossing
+Industrial & Scientific MovieLens-1M Book Crossing
 Recall NDCG HR Recall NDCG HR Recall NDCG HR
@@
 Imp.over the best zero-shot baseline $157.09\%$ $127.69\%$ $30.29\%$ $66.67\%$ $64.16\%$ $37.78\%$ $101.55\%$ $63.71\%$ $47.97\%$
-Results.
-Table 4 presents the zero-shot recommendation performance comparison on entirely new datasets.
-The best-performing methods are bold and starred, while the second-best methods are underlined.
-We observe that:
+Findings. Table 4 presents the zero-shot recommendation performance comparison. The best methods are bold and starred, while the second-best methods are underlined. We observe that:
+  • Advanced language representations provide opportunities for learning transferable item representations.
 AlphaRec demonstrates strong zero-shot recommendation capabilities, comparable to or even surpassing the fully trained LightGCN.
-On datasets from completely different platforms (e.g., MovieLens-1M and Book Crossing), AlphaRec is comparable with the fully trained LightGCN.
-On the same Amazon platform dataset, Industrial, AlphaRec even surpasses LightGCN, for which we attribute to the possibility that AlphaRec implicitly learns unique user behavioral patterns on the Amazon platform [1].
+AlphaRec performs better on the Amazon Industrial & Scientific dataset, possibly because it captures user behavioral patterns of the same platform (Ni et al., 2019) through training on multiple Amazon datasets.
 Conversely, ZESRec and UniSRec exhibit a marked performance decrement compared with AlphaRec.
 We attribute this phenomenon to two aspects.
-On the one hand, BERT-style LMs [4, 5] used in these works may not have effectively encoded user preference similarities, which is consistent with our previous findings in Section 2.
-On the other hand, components designed for the next item prediction task in sequential recommendation [86] may not be suitable for capturing the general preferences of users in CF scenarios.
-  • The zero-shot recommendation capability of AlphaRec generally benefits from an increased amount of training data, without harming the performance on source datasets.
-As illustrated in Figure 8, the zero-shot performance of AlphaRec, when trained on a mixed dataset, is generally superior to training on one single dataset [37].
-Additionally, we also note that training data with themes similar to the target domain contributes more to the zero-shot performance.
-For instance, the zero-shot capability on MovieLens-1M may primarily stem from Movies & TV.
-Furthermore, we discover that AlphaRec, when trained jointly on multiple datasets, hardly experiences a performance decline on each source dataset.
-These findings further point to the general recommendation capability of a single pre-trained AlphaRec across multiple datasets.
+On the one hand, BERT-style LMs (Devlin et al., 2019; Liu et al., 2019) used in these works may not have effectively encoded user preference similarities, which is consistent with our previous findings in Section 3.
+On the other hand, components designed for the next item prediction task in sequential recommendation (Kang & McAuley, 2018) may not be suitable for capturing the general preferences of users in CF scenarios.
+Moreover, AlphaRec also outperforms the leading LLM-based zero-shot recommender LLMRank (see Appendix D.2.3).
+  • The zero-shot recommendation capability of advanced language representations generally benefits from an increased amount of training data, without compromising performance on source datasets.
+As illustrated in Table 9, the zero-shot performance of AlphaRec, when trained on a mixed dataset, is generally superior to training on one single dataset (Hou et al., 2022).
+Moreover, we discover that AlphaRec, when trained jointly on multiple datasets, hardly experiences a performance decline on each source dataset.
+These results indicate the general recommendation capability of a single pre-trained AlphaRec across multiple datasets.
 The above findings also offer a potential research path to achieve general recommendation capabilities, by incorporating more training data with more themes.
-See more details about these results in Appendix C.3.3.
-4.3 User Intention Capture Performance (RQ3)
-Motivation.
-We aim to investigate whether a straightforward paradigm shift enables pre-trained AlphaRec to perceive text-based user intentions and refine recommendations.
-Specifically, we introduce an adjustable hyperparameter $\alpha$ to combine user intentions with historical interests.
-Task and datasets.
+See more details about these results in Appendix D.2.4.
+5.3 Intention-aware Ability (Potential 3)
+The language understanding ability in advanced language representations (especially representations from LLM-based text embedding models) offers the opportunity for perceiving text-based user intentions and refining recommendations.
+To study the potential of intention-aware ability, we introduce a new hyperparameter $\alpha$ in AlphaRec to combine user intentions with historical interests.
+Experimental settings.
+To endow AlphaRec with user intention-aware ability, we adopt a simple paradigm shift by introducing a user intention representation $\mathbf{e}_{u}^{Intention}$.
+In the inference stage, we obtain the language representation $\mathbf{e}_{u}^{Intention}$ for each user intention query and combine it with the original user representation to get a new user representation as $\mathbf{\tilde{e}}_{u}^{(0)}=(1-\alpha)\mathbf{e}_{u}^{(0)}+\alpha\mathbf{e}_{u}^{Intention}$ (Ai et al., 2017).
+This new user representation $\mathbf{\tilde{e}}_{u}^{(0)}$ is sent into the pre-trained AlphaRec for recommendation.
 We test the user intention capture ability of AlphaRec on MovieLens-1M and Video Games.
-In the test set, only one target item remains for each user [87], with one intention query generated by ChatGPT [88, 40] (see the details about how to generate and check these intention queries in Appendix C.4.1).
-In the training stage, we follow the same procedure as illustrated in Section 2 to train AlphaRec.
-In the inference stage, we obtain the language representation $\mathbf{e}_{u}^{Intention}$ for each user intention query and combine it with the original user representation to get a new user representation as $\mathbf{\tilde{e}}_{u}^{(0)}=(1-\alpha)\mathbf{e}_{u}^{(0)}+\alpha\mathbf{e}_{u}^{Intention}$ [87].
-This new user representation is sent into the freezed AlphaRec for recommendation.
+In the test set, only one target item remains for each user (Ai et al., 2017), with one intention query generated by ChatGPT (OpenAI, 2023; Hou et al., 2024a) (see the details about how to generate and check these intention queries in Appendix D.3.1).
 We report a relatively small $K$ = 5 for all metrics to better reflect the intention capture accuracy.
+Findings: We report the user intention capture experiment results in Table 5, show one case study in Figure 4(a), and study the effect of $\alpha$ in Figure 4(b). We find out that:
+The language understanding ability in advanced language representations enables recommenders to perceive user intentions and refine recommendations.
+As shown in Table 5, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention).
+Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020) by a large margin, even without additional training on search tasks.
+We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more case study results in Appendix D.3.3).
+Additionally, the intention-aware ability also benefits from user historical interests.
+Figure 4(b) depicts the effect of $\alpha$.
+The $\alpha$ controls the strength of user intention, where $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored.
+The convex curve in Figure 4(b) suggests that both user historical interests and user intention play vital roles.
+The above findings suggest the potential of adopting advanced language representations to perceive text-based user intentions and refining recommendations (see more details in Appendix D.3).
 Table 5: The performance comparison in user intention capture.
@@
 Figure 4: User intention capture experiments on MovieLens-1M.
-(3(a)) AlphaRec refines the recommendations according to language-based user intention.
-(3(b)) The effect of user intention strength $\alpha$.
-User intention capture results.
-Table 5 represents the user intention capture experiment results, compared with the baseline TEM [89].
-Clearly, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention).
-Moreover, AlphaRec outperforms the baseline model TEM by a large margin, even without additional training on search tasks.
-We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user (see more case study results in Appendix C.4.3).
-As shown in Figure 3(a), AlphaRec accurately captures the hidden user intention for “Godfather”, while keeping most of the recommendation results unchanged.
-This indicates that AlphaRec captures the user intention and historical interests simultaneously.
-Effect of the intention strength $\alpha$.
-By controlling the value of $\alpha$, AlphaRec can provide better recommendation results, with a balance between user historical interests and user intent capture.
-Figure 3(b) depicts the effect of $\alpha$.
-Initially, as $\alpha$ increases, the recommendation performance rises accordingly, indicating that incorporating user intention enables AlphaRec to provide better recommendation results.
-However, as the $\alpha$ approaches 1, the recommendation performance starts to decrease, which suggests that the user historical interests learned by AlphaRec also play a vital role.
-The similar effect of $\alpha$ on Video Games is discussed in Appendix C.4.4.
+(4(a)) AlphaRec refines the recommendations according to language-based user intention.
+(4(b)) The effect of user intention strength $\alpha$.
 6 Limitations
-There are several limitations not addressed in this paper.
-On the one hand, although we have demonstrated the excellence of AlphaRec for multiple tasks on various offline datasets, the effectiveness of online employment remains unclear.
-On the other hand, although we have successfully explored the potential of language-representation-based recommenders by incorporating essential components in leading CF models, we do not elaboratively focus on designing new components for CF models.
+There are several limitations unaddressed in this paper.
+On the one hand, while we explore the relationship between language spaces and behavior spaces through empirical results, there is no theoretical guarantee.
+On the other hand, although we investigate the potential of advanced language representations for recommendation, we do not design any new components for CF models.
 7 Conclusion
-In this paper, we explored what knowledge about recommendations has been encoded in the LM representation space.
-Specifically, we found that the advanced LMs representation space exhibits a homomorphic relationship with an effective recommendation space.
-Based on this finding, we developed a simple yet effective CF model called AlphaRec, which exhibits good recommendation performance with zero-shot recommendation and user intent capture ability.
-We pointed out that AlphaRec follows a new recommendation paradigm, language-representation-based recommendation, which uses language representations from LMs to represent users and items and completely abandons ID-based embeddings.
-We believed that AlphaRec is an important stepping stone towards building general recommenders in the future.
-[^1] Thc broader impact of AlphaRec will be detailed in Appendix E
+In this paper, we explored the relationship between language space and behavior spaces for recommendation, and explored the potential for using language representations for recommendation.
+Empirical results suggest the possible presence of homomorphism between advanced LMs representation spaces and an effective item representation space for recommendation.
+Inspired by this finding, we discussed how to unleash the power of advanced language representations by developing a simple yet effective CF model called AlphaRec.
+Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: fast convergence, zero-shot ability, and intention-aware ability.
+Possible future work will involve exploring the space relationship from both theoretical and multimodal perspectives.
+We believed that this paper sheds light on rethinking the connection between language modeling and user behavior modeling, benefiting both natural language processing and recommender system communities.
 Appendix A Related Works
@@
 The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space.
-An important and effective approach for interpreting and analyzing representations of LMs is linear probing [68, 67].
-The main idea of linear probing is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure [9] ) from the representations in the hidden layers of LMs.
-A high probing result (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations [68, 10].
-Recent studies empirically demonstrate that concepts such as color [90], game states [8].
+Linear methods (e.g., linear probing (Merullo et al., 2023) and linear mapping (Alain & Bengio, 2017)) are important and effective approaches for interpreting and analyzing representations of LMs (Ravichander et al., 2021).
+The main idea of linear methods is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure (Vulic et al., 2020) ) from the representations in the hidden layers of LMs, or transforming language representations into another feature space with a linear matrix.
+A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023).
+Recent studies empirically demonstrate that concepts such as color (Patel & Pavlick, 2022), game states (Li et al., 2023a).
 and geographic position are encoded in LMs.
-Furthermore, these concepts may even be linearly encoded in the representation space of LMs [8, 91].
+Furthermore, these concepts may even be linearly encoded in the representation space of LMs (Li et al., 2023a; Park et al., 2023).
 Collaborative filtering.
-Collaborative filtering (CF) [92] is an advanced technique in modern recommender systems.
-The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup [55].
-Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern [71].
-Recent advances in CF mainly benefit from two aspects, graph convolution [75] and contrastive learning [92].
-These CF models exhibit superior recommendation performance by conducting the embedding propagation [52, 56] and applying contrastive learning objectives [83, 63, 70].
+Collaborative filtering (CF) (Ren et al., 2024a) is an advanced technique in modern recommender systems.
+The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup (Koren et al., 2009).
+Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern (Rendle et al., 2012).
+Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a).
+These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024).
 However, although effective, these methods are still limited, due to the ID-based paradigm.
-Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains [37] or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) [34].
+Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains (Hou et al., 2022) or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) (Yuan et al., 2023).
 LMs for recommendation.
 The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation.
-The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LLM-based recommendation.
-The first research direction, LLM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs [48, 47, 46, 49, 93, 94].
+The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LM-as recommender.
+The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024c).
 Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences.
-The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) [34, 35].
+The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) (Yuan et al., 2023; Li et al., 2023c).
 These methods tend to train the LM as the text modality encoder together with the traditional recommender.
 In previous studies, BERT-style LMs are widely used as the text modality encoder.
-The third research line, LLM-based recommendation, directly uses LLMs as the recommender and recommends items in a text generation paradigm.
-Early attempts focus on adopting in-context learning (ICL) [95] and prompting pre-trained LLMs [96, 97, 98, 99].
+The third research line fails in directly using LMs as the recommender and recommends items in a text generation paradigm.
+Early attempts focus on adopting in-context learning (ICL) (Dong et al., 2022) and prompting pre-trained LMs (Hou et al., 2024b; Liu et al., 2023a; Dai et al., 2023; Gao et al., 2023).
 However, such naive methods tend to yield poor performance compared to traditional models.
-Therefore, recent studies concentrate on fine-tuning LLMs on recommendation-related corpus [16, 15, 26, 25, 29] and align the LLMs with the representations from traditional recommenders as the additional modality [17, 20, 27, 100].
-Appendix B Linear Mapping
+Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023d; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024b; Hua et al., 2023; Chen et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).
+Appendix B Uncovering Collaborative Signals in LMs via Linear Mapping
 B.1 Brief of Used LMs
-We briefly introduce the LMs we use for linear mapping in Section 2.
-  • BERT [4] is an encoder-only language model based on the transformer architecture [3], pre-trained on text corpus with unsupervised tasks.
+We briefly introduce the LMs we use for linear mapping in Section 3.1.
+  • BERT (Devlin et al., 2019) is an encoder-only language model based on the transformer architecture (Vaswani et al., 2017), pre-trained on text corpus with unsupervised tasks.
 BERT adopts bidirectional self-attention heads to learn bidirectional representations.
-  • RoBERTa [5] is an enhanced version of BERT.
+  • RoBERTa (Liu et al., 2019) is an enhanced version of BERT.
 RoBERTa preserves the architecture of BERT but improves it by training with more data and large batches, adopting dynamic masking, and removing the next sentence prediction objective.
-  • Llama2-7B [6] is an open-source decoder-only LLM with 7 billion parameters.
-Llama2 adopts grouped-query attention, with longer context length and larger size of the pre-training corpus compared with Llama-7B [101].
-  • Mistral-7B [72] is an open-source pre-trained decoder-only LLM with 7 billion parameters.
+  • Llama2-7B (Touvron et al., 2023b) is an open-source decoder-only LLM with 7 billion parameters.
+Llama2 adopts grouped-query attention, with longer context length and larger size of the pre-training corpus compared with Llama-7B (Touvron et al., 2023a).
+  • Mistral-7B (Jiang et al., 2023) is an open-source pre-trained decoder-only LLM with 7 billion parameters.
 Mistral 7B leverages grouped-query attention, coupled with sliding window attention for faster and lower cost inference.
-  • text-embedding-ada-v2 & text-embeddings-3-large [73] are leading text embedding models released by OpenAI.
+  • text-embedding-ada-v2 & text-embeddings-3-large (Neelakantan et al., 2022) are leading text embedding models released by OpenAI.
 These models are built upon decoder-only GPT models, pre-trained on unsupervised data at scale.
-  • SFR-Embedding-Mistral [74] is a decoder-based text embedding model built upon the open-source LLM Mixtral-7B [72].
+  • SFR-Embedding-Mistral (Meng et al., 2024) is a decoder-based text embedding model built upon the open-source LLM Mixtral-7B (Jiang et al., 2023).
 SFR-Embedding-Mistral introduces task-homogeneous batching and computes contrastive loss on “hard negatives”, which brings a better performance than the vanilla Mixtral-7B model.
-B.2 Extracting Representations from LMs
+B.2 Generating Item Representations from LMs
 We present how to extract representations from LMs.
-For encoder-based LMs (e.g., BERT [4] and RoBERTa [5]), we use the representation of the last hidden state corresponding to the [CLS] token [40].
-For decoder-based models (e.g., Llama-7B [6, 72], Mistral-7B, and SFR-Embedding-Mistral [74]), we use the representation in the last transformer block [3], corresponding to the last input token [10, 102, 73].
-Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^2] https://platform.openai.com/docs/guides/embeddings [73]), we directly call the API interface to obtain representations.
-B.3 Empirical Findings
-We examinate whether the better performance of leading LMs comes from the more compact representation ability.
-We randomly shuffle item representations and conduct the same linear mapping experiment.
-As illustrated in Table 6, randomly shuffled representations, text-embeddings-3-large (Random), lags largely behind the vanilla linear mapping method, suggesting that the performance improvement not merely comes from the better feature encoding ability.
-Table 6: Linear mapping performance of randomly shuffled item representations
-Books Movies & TV Video Games
-Recall NDCG HR Recall NDCG HR Recall NDCG HR
-BERT 0.0226 0.0194 0.1240 0.0415 0.0399 0.2362 0.0524 0.0309 0.1245
-text-embeddings-3-large (Random) 0.0200 0.0197 0.1316 0.0559 0.0528 0.3204 0.0562 0.0328 0.1351
-text-embeddings-3-large 0.0735 0.0608 0.3355 0.1109 0.1023 0.5200 0.1367 0.0793 0.2928
-Appendix C Experiments
+For encoder-based LMs (e.g., BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019)), we use the representation of the last hidden state corresponding to the [CLS] token (Hou et al., 2024a).
+For decoder-based models (e.g., Llama-7B (Touvron et al., 2023b; Jiang et al., 2023), Mistral-7B, and SFR-Embedding-Mistral (Meng et al., 2024)), we use the representation in the last transformer block (Vaswani et al., 2017), corresponding to the last input token (Gurnee & Tegmark, 2023; Todd et al., 2023; Neelakantan et al., 2022).
+Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^2] https://platform.openai.com/docs/guides/embeddings (Neelakantan et al., 2022)), we directly call the API interface to obtain representations.
+Appendix C Leveraging Language Representations for Better Recommendation
 C.1 Datasets
 Table 6: Dataset statistics.
-Books Movies & TV Video Games Industrial MovieLens-1M Book Crossing
+Books Movies & TV Video Games Industrial & Scientific MovieLens-1M Book Crossing
 #Users 71,306 26,073 40,834 15,141 6,040 6,273
@@
 Density 0.0008 0.0026 0.0007 0.0010 0.0542 0.0076
-We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com [1] (i.e., Books, Movies & TV, Video Games, and Industrial), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing).
-Table 7 reports the data statistics of each dataset.
-We divide the history interaction of each user into training, validation, and testing sets with a ratio of 4:3:3, and remove users with less than 20 interactions following previous studies [50].
+We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing).
+Table 6 reports the dataset statistics.
+We divide the history interaction of each user into training, validation, and testing sets with a ratio of 4:3:3, and remove users with less than 20 interactions following previous studies (Zhang et al., 2023b).
 We also remove items from the test and validation sets that do not appear in the training set, to address the cold start problem.
@@
 Item Title Examples Books: Dismissed with Prejudice: A J.P.
-Beaumont Novel; Die for Love: A Jacqueline Kirby Novel of Suspense; The Cloud; Memories Before and After the Sound of Music: An Autobiography; Harry Potter and the Sorcerer’s Stone; Movies & TV: Batman Begins; Fantastic Four; Max Headroom: The Complete Series; Madagascar; Land of the Dead; King Kong; Video Games: Fighting Force; Tomb Raider II; Tomb Raider; WWF Warzone; Kartia: The Word of Fate; Snowboard Kids; Command & amp; Conquer: Tiberian Sun - PC; Final Fantasy VII; Grim Fandango - PC; Half-Life - PC; MovieLens-1M: Basquiat (1996); Tin Cup (1996); Godfather, The (1972); Supercop (1992); Manny & Lo (1996); Bound (1996); Carpool (1996); Book Crossing: Prague : A Novel; Chocolate Jesus; Wie Barney es sieht; To Kill a Mockingbird; Sturmzeit.
-Roman; A Soldier of the Great War; Pride and Prejudice (Dover Thrift Editions); Industrial: Jurassic Perisphinctes Ammonites from France; FS9140: Spinosaurus - Dinosaur Tooth 20-30mm; FS9410: USA Eocene, Fossil Fish (Knightia alt), A-grade; Delta 50-857 Charcoal Filter for 50-868; Hitachi RP30SA 7-1/2 Gallon Stainless Steel Industrial Shop Vacuum (Discontinued by Manufacturer); Makita 632002-4 14-Inch Cut-Off Wheels (5-Pack) (Discontinued by Manufacturer); PORTER-CABLE 740001801 4 1/2-Inch by 10yd 180 Grit Adhesive-Backed Sanding Roll;
+Beaumont Novel; Die for Love: A Jacqueline Kirby Novel of Suspense; The Cloud; Memories Before and After the Sound of Music: An Autobiography; Harry Potter and the Sorcerer’s Stone; Movies & TV: Batman Begins; Fantastic Four; Max Headroom: The Complete Series; Madagascar; Land of the Dead; King Kong; Video Games: USB Microphone for RockBand or Guitar Hero (PS3, Wii, Xbox360); Command &amp; Conquer: Tiberian Sun - PC; Tomb Raider III: Adventures of Lara Croft; Kartia: The Word of Fate; Snowboard Kids; Command & amp; Conquer: Tiberian Sun - PC; Final Fantasy VII; Grim Fandango - PC; Half-Life - PC; MovieLens-1M: Basquiat (1996); Tin Cup (1996); Godfather, The (1972); Supercop (1992); Manny & Lo (1996); Bound (1996); Carpool (1996); Book Crossing: Prague : A Novel; Chocolate Jesus; Wie Barney es sieht; To Kill a Mockingbird; Sturmzeit.
+Roman; A Soldier of the Great War; Pride and Prejudice (Dover Thrift Editions); Industrial & Scientific: Jurassic Perisphinctes Ammonites from France; FS9140: Spinosaurus - Dinosaur Tooth 20-30mm; FS9410: USA Eocene, Fossil Fish (Knightia alt), A-grade; Delta 50-857 Charcoal Filter for 50-868; Hitachi RP30SA 7-1/2 Gallon Stainless Steel Industrial Shop Vacuum (Discontinued by Manufacturer); Makita 632002-4 14-Inch Cut-Off Wheels (5-Pack) (Discontinued by Manufacturer); PORTER-CABLE 740001801 4 1/2-Inch by 10yd 180 Grit Adhesive-Backed Sanding Roll;
 Figure 5: Example of item titles.
-C.2 General Recommendation
-C.2.1 Baselines
-We incorporate a series of CF models as our baselines for general recommendation.
+C.2 Baselines
+We incorporate a series of ID-based CF models as our baselines for general recommendation.
 These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec).
-For these LM-enhanced CF methods, we adopt the leading CF method XSimGCL as the backbone.
-  • MF [55, 71] is the most basic CF model. It denotes users and items with ID-based embeddings and conducts matrix factorization with Bayesian personalized ranking (BPR) loss.
-  • MultVAE [64] is a traditional CF model based on the variational autoencoder (VAE).
+For these LM-enhanced CF methods, we adopt the leading method XSimGCL as the backbone.
+  • MF (Koren et al., 2009; Rendle et al., 2012) is the most basic CF model.
+It denotes users and items with ID-based embeddings and conducts matrix factorization with Bayesian personalized ranking (BPR) loss.
+  • MultVAE (Liang et al., 2018) is a traditional CF model based on the variational autoencoder (VAE).
 It regards the item recommendation as a generative process from a multinomial distribution and uses variational inference to estimate parameters.
 We adopt the same model structure as suggested in the paper: 600 → 200 → 600.
-  • LightGCN [56] is a light graph convolution network tailored for the recommendation, which deletes redundant feature transformation and activation function in NGCF [52].
-  • SGL [83] introduces graph contrastive learning into recommender models for the first time.
+  • LightGCN (He et al., 2021) is a light graph convolution network tailored for the recommendation, which deletes redundant feature transformation and activation function in NGCF (Wang et al., 2019b).
+  • SGL (Wu et al., 2021) introduces graph contrastive learning into recommender models for the first time.
 By employing node or edge dropout to generate augmented graph views and conduct contrastive learning between two views, SGL achieves better performance than LightGCN.
-  • BC Loss [79] introduces a robust and model-agnostic contrastive loss, handling various data biases in recommendation, especially for popularity bias.
-  • XSimGCL [70] directly generates augmented views by adding noise into the inner layer of LightGCN without graph augmentation.
+  • BC Loss (Zhang et al., 2022) introduces a robust and model-agnostic contrastive loss, handling various data biases in recommendation, especially for popularity bias.
+  • XSimGCL (Yu et al., 2024) directly generates augmented views by adding noise into the inner layer of LightGCN without graph augmentation.
 The simplicity of XSimGCL leads to a faster convergence speed and better performance.
-  • KAR [48] enhances recommender models by integrating knowledge from large language models (LLMs).
+  • KAR (Xi et al., 2023) enhances recommender models by integrating knowledge from LMs.
 It generates textual descriptions of users and items and combine the LM representations with traditional recommenders using a hybrid-expert adaptor.
-  • RLMRec [47] aligns semantic representations of users and items with the representations in CF models through a contrastive loss, as an additional loss trained together with the CF model.
+  • RLMRec (Ren et al., 2024b) aligns semantic representations of users and items with the representations in CF models through a contrastive loss, as an additional loss trained together with the CF model.
 The fusion of semantic information and collaborative information brings performance improvement.
-C.2.2 Ablation Study
+C.3 Ablation Study
 (a) Ablation study on Movies & TV
@@
 Figure 6: Ablation study
-We conduct the same ablation study as introduced in Section 4.1 on Movies & TV and Video Games datasets.
-As illustrated in Figure 5, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.1.
-C.2.3 The t-SNE Visualization Comparison
+We conduct the same ablation study as introduced in Section 4 on Movies & TV and Video Games datasets.
+As illustrated in Figure 6, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.
+C.4 The t-SNE Visualization Comparison
 (a) LM representations
@@
 Figure 7: The t-SNE visualization of representations on Movies & TV.
-(6(a)) The item representations in the LM space.
-(6(b)) The item representations obtained by replacing the MLP with a linear mapping matrix in AlphaRec.
-(6(c)) The item representations obtained from AlphaRec.
+(7(a)) The item representations in the LM space.
+(7(b)) The item representations obtained by replacing the MLP with a linear mapping matrix in AlphaRec.
+(7(c)) The item representations obtained from AlphaRec.
 In this section, we aim to intuitively explore how the MLP in AlphaRec further helps in excavating collaborative signals in language representations, compared to the linear mapping matrix.
-We visualize the item representations from LMs, AlphaRec (w/o MLP), and AlphaRec in Figure 6, where AlphaRec (w/o MLP) denotes replacing the MLP with a linear mapping matrix.
+We visualize the item representations from LMs, post-mapping representations from AlphaRec (w/o MLP), and post-mapping representations from AlphaRec in Figure 7, where AlphaRec (w/o MLP) denotes replacing the MLP with a linear mapping matrix.
 We observed that movies about superhero and monster cluster in all representation spaces, indicating both AlphaRec (w/o MLP) and AlphaRec capture the preference similarities between these items and preserve the clustering relationship.
@@
 These results indicate that AlphaRec exhibits a more fine-grained preference capture ability with the help of nonlinear transformation.
-C.3 Zero-shot Recommendation
-C.3.1 Co-training on Multiple Datasets
+Appendix D Exploring Potentials of Language Representations for Recommendation
+D.1 Fast Convergence Speed
+We report the training cost of AlphaRec in this section.
+Table 7 reports the seconds needed per epoch and the total training cost until convergence.
+Here Amazon-Mix denotes the mixed dataset of Books, Movies & TV, and Video Games.
+It’s worth noting that AlphaRec converges quickly and only requires a small amount of training time.
+Table 7: Training cost of AlphaRec (seconds per epoch/in total).
+Books Movies & TV Video Games Amazon-Mix
+AlphaRec 40.1 / 1363.4 12.3 / 479.7 7.4 / 214.6 107.2 / 5788.8
+D.2 Zero-shot Ability
+D.2.1 Co-training on Multiple Datasets
 Co-training on multiple datasets is similar to training on one single dataset, where the only difference lies in the negative sampling.
@@
 The other training procedures remain the same with training on one single dataset.
-C.3.2 Baselines
-Since previous works about zero-shot recommendation mostly focus on sequential recommendation [86, 85], we slightly modify two methods in sequential recommendation, ZESRec [38] and UniSRec [37] as our baselines.
+D.2.2 Baselines
+Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify two methods in sequential recommendation, ZESRec (Ding et al., 2021) and UniSRec (Hou et al., 2022) as our baselines.
 Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
@@
   • Pop denotes randomly recommending from the most popular items. Here popularity denotes the number of users that have interacted with the item.
-  • ZESRec [38] is the first work that defines the problem of zero-shot recommendation.
+  • ZESRec (Ding et al., 2021) is the first work that defines the problem of zero-shot recommendation.
 To address this problem, this work introduces a hierarchical Bayesian model with representations from the pre-trained BERT.
-  • UniSRec [37] aims to learn universal item representations from BERT, with parametric whitening and a MoE-enhanced adaptor.
+  • UniSRec (Hou et al., 2022) aims to learn universal item representations from BERT, with parametric whitening and a MoE-enhanced adaptor.
 By pre-training on multiple source datasets, UniSRec can conduct zero-shot recommendation on various datasets in a transductive or inductive paradigm.
-C.3.3 The Effect of Training Datasets
+D.2.3 Comparison with LLMRank
+Table 8: Zero-shot performance comparison with LLMRank
+MovieLens-1M Steam
+NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG@1 NDCG@5 NDCG@10 NDCG@20
+LLMRank 0.2485 0.4115 0.5249 0.5612 0.3112 0.4413 0.5255 0.5302
+AlphaRec 0.3919 0.6038 0.6543 0.6672 0.4450 0.6131 0.6394 0.6714
+Imp .57.71% 46.73% 24.65% 18.89% 42.99% 38.93% 21.67% 26.63%
+Table 8 illustrates the zero-shot recommendation performance compared with the LLM4Rec method LLMRank.
+We adopt the same setting of LLMRank, equipping 19 negative items for each positive item, and evaluate the NDCG on the candidate set.
+AlphaRec exhibits excellent zero-shot performance, significantly surpassing LLMRank.
+Moreover, the improvement over LLMRank exhibits a rising trend as the $K$ of NDCG decreases.
+D.2.4 The Effect of Training Datasets
 The effect of the training dataset scale on zero-shot recommendation.
@@
 Table 9: The effect of the training dataset on zero-shot recommendation
-Industrial MovieLens-1M Book Crossing
+Industrial & Scientific MovieLens-1M Book Crossing
 Recall NDCG HR Recall NDCG HR Recall NDCG HR
@@
 AlphaRec (trained on mixed dataset) 0.0979 0.0818 0.4147 0.1194 0.1107 0.5463 0.1381 0.0827 0.2985
-C.4 User Intention Capture
-C.4.1 Intention Query Generation
+D.3 Intention-aware Ability
+D.3.1 Intention Query Generation
 Intention Query Generation Input
@@
 For each item in the dataset, we generate a fixed user intention query.
-Following the previous work [40], we generate user intention queries with the help of ChatGPT [88].
-As shown in Figure 7, we prompt ChatGPT in a Chain-of-Thought (CoT) [103] paradigm and adopt the output as the user intention query.
+Following the previous work (Hou et al., 2024a), we generate user intention queries with the help of ChatGPT (OpenAI, 2023).
+As shown in Figure 8, we prompt ChatGPT in a Chain-of-Thought (CoT) (Wei et al., 2022) paradigm and adopt the output as the user intention query.
 We adopt a rule-based strategy to ensure the quality of generated queries, and regenerate the wrong query.
 Considering the huge amount of item title text, we use ChatGPT3.5 API for generating all queries for the budget’s sake.
-C.4.2 Baseline
-AlphaRec exhibits user intention capture abilities, although not specially designed for search tasks. We compare AlphaRec with TEM [89] which falls in the field of personalized search [87, 104].
-  • TEM [89] uses a transformer to encode the intention query together with user history behaviors, which enables it to achieve better search results by considering the user’s historical interest.
-C.4.3 Case Study
+D.3.2 Baseline
+AlphaRec exhibits user intention capture abilities, although not specially designed for search tasks.
+We compare AlphaRec with TEM (Bi et al., 2020) which falls in the field of personalized search (Ai et al., 2017; McAuley et al., 2015).
+  • TEM (Bi et al., 2020) uses a transformer to encode the intention query together with user history behaviors, which enables it to achieve better search results by considering the user’s historical interest.
+D.3.3 Case Study
 We conduct two more case studies to verify the user intention capture ability of AlphaRec.
@@
 Figure 10: Case study of user intention capture on Video Games
-C.4.4 Effect of the Intention Strength Alpha
+D.3.4 Effect of the Intention Strength Alpha
 The value of $\alpha$ controls the balance between the user’s historical interests and the user intention query.
@@
 Figure 11: Effect of $\alpha$ on Video Games
-C.5 Trainig Cost
-We report the training cost of AlphaRec in this section.
-Table 10 reports the seconds needed per epoch and the total training cost until convergence.
-Here Amazon-Mix denotes the mixed dataset of Books, Movies & TV, and Video Games.
-It’s worth noting that AlphaRec converges quickly and only requires a small amount of training time.
-Table 10: Training cost of AlphaRec (seconds per epoch/in total).
-Books Movies & TV Video Games Amazon-Mix
-AlphaRec 40.1 / 1363.4 12.3 / 479.7 7.4 / 214.6 107.2 / 5788.8
-Appendix D Hyperparameter Settings and Implementation Details
+Appendix E Hyperparameter Settings and Implementation Details
 Table 11: Hyperparameters search space for baselines.
@@
 In AlphaRec, the dimensions of the input and output in the two-layer MLP are 3072 and 64 respectively, with the hidden layer dimension as 1536.
-We apply the all-ranking strategy [105] for all experiments, which ranks all items except positive ones in the training set for each user.
+We apply the all-ranking strategy (Krichene & Rendle, 2020) for all experiments, which ranks all items except positive ones in the training set for each user.
 We search hyperparameters for baselines according to the suggestion in the literature.
@@
 We adopt $\tau$ = 0.2 for the MovieLens-1M dataset for the user intention capture experiment in Section 5.3.
-Appendix E Broader Impact
-The proposed AlphaRec can significantly improve the performance of zero-shot recommendation and the capability of user intent capture, offering a good approach to crafting more personalized recommendation results.
-One concern of AlphaRec is the potential for the representations generated by language models can be maliciously attacked, which may result in erroneous or unexpected recommendations.
-Therefore, we kindly advise researchers to cautiously check the quality of the language representations before using AlphaRec.

### v3

@@
-Do Language Models Encode Collaborative Signals in Recommendation? Findings and Potentials
+Language Representations Can be What Recommenders Need: Findings and Potentials
 Abstract
@@
 Moreover, we systematically analyze this simple model and find several key features for using advanced language representations: a good initialization for item representations, superior zero-shot recommendation abilities in new datasets, and being aware of user intention.
-Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities  [^1] Codes are available at https://github.com/LehengTHU/AlphaRec.
+Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities.
 1 Introduction
@@
 Specifically, on the one hand, when using LMs as recommenders to directly output items of interest, aligning the language space with the behavior space can significantly improve the recommendation performance (Lin et al., 2023a; Vats et al., 2024; Xu et al., 2024).
-Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023d; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024).
-On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024a), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; c).
+Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023e; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023f; Yang et al., 2023; Tennenholtz et al., 2024), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024).
+On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024b), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; d).
 Despite these efforts, explicit studies of the relationship between language and behavior spaces remain largely unexplored in the recommendation domain.
@@
 LM-based generator.
-Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024b; Zhang et al., 2024a; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$.
+Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024c; Zhang et al., 2024b; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$.
 These features are fed into the LM-based generator, typically a combination of two subsequent components: (1) A frozen LM to extract $i$’s language representation $\mathbf{z}_{i}$, such as the encoder-only LMs like BERT-style models (Devlin et al., 2019; Liu et al., 2019), the decoder-only LMs like Llama-style autoregressive models (Touvron et al., 2023b; Jiang et al., 2023) and OpenAI text embedding models (Neelakantan et al., 2022); and (2) A trainable projector to map $\mathbf{z}_{i}$ into the final representation $\mathbf{e}_{i}$, often using layers like graph convolution layers (He et al., 2021).
@@
 We compare AlphaRec with leading ID-based CF baselines, to assess the effectiveness of adopting advanced language representations.
-We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023).
 In addition to classic baselines introduced in section 3.2, we consider two categories of leading ID-based CF baselines, CL-based CF methods: SGL (Wu et al., 2021), BC Loss (Zhang et al., 2022), XSimGCL (Yu et al., 2024) and LM-enhanced CF methods: KAR (Xi et al., 2023), RLMRec (Ren et al., 2024b).
-See more details about baselines in Appendix C.2.
-Recommendation capabilities (RQ3). Table 3 presents performance comparison. The best-performing methods are bold, while the second-best methods are underlined. We observe that:
+We also consider a variant of XSimGCL (i.e., XSimGCLt) where ID embeddings are replaced with language representations, and directly using language representations $\mathbf{z}_{u}$ and $\mathbf{z}_{i}$ for recommendation without training (i.e., EMB-KNN).
+See baseline details in Appendix C.2.
+Recommendation capabilities (RQ3).
+Table 3 presents performance comparison.
+To alleviate the possible information leakage problem that pretraining data of LMs may include the training data, we also conduct experiments on a newly published dataset Amazon 2023 in Appendix C.6.
+The best-performing methods are bold, while the second-best methods are underlined.
+We observe that:
 Table 3: The performance comparison with ID-based CF baselines. The improvement achieved by AlphaRec is significant ($p$-value <<much-less-than$<$ < 0.05).
@@
 XSimGCL (Yu et al., 2024) 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550 0.0879 0.0745 0.3918
+XSimGCLt (Yu et al., 2024) 0.1015 0.0951 0.5016 0.1199 0.0679 0.2674 0.0900 0.0736 0.4036
 KAR (Xi et al., 2023) 0.1084 0.1001 0.5134 0.1181 0.0693 0.2571 0.0852 0.0734 0.3834
 RLMRec (Ren et al., 2024b) 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997 0.0928 0.0774 0.4092
+EMB-KNN 0.0548 0.0380 0.2916 0.0879 0.0389 0.1970 0.0434 0.0248 0.1851
 AlphaRec 0.1221* 0.1144* 0.5587* 0.1519* 0.0894* 0.3207* 0.0991* 0.0828* 0.4185*
@@
 AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets, with an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline.
-Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results in Appendix C.3) in unleashing the power of language representations.
-Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively.
-Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance.
-These findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
+Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results and details in Appendix C.3) in unleashing the power of language representations.
+Moreover, simply replacing the ID embeddings in XSimGCL with language representations (i.e., XSimGCLt) does not bring a stable and remarkable improvement, which is consistent with previous findings (Yuan et al., 2023; Li et al., 2023c).
+This phenomenon highlights the importance of appropriate model design beyond simple input feature replacement.
+Above findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
   • The incorporation of advanced language representations can benefit traditional ID-based CF methods.
@@
 Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec.
-We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024c).
+We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024d).
 5 Exploring Potentials of Language Representations for Recommendation
@@
 Experimental settings.
-In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024b), which is different from the research line of cross-domain recommendation (Zhu et al., 2021).
+In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024c), which is different from the research line of cross-domain recommendation (Zhu et al., 2021).
 We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M (Harper & Konstan, 2016), Book Crossing (Lee et al., 2019), and Amazon Industrial & Scientific (Ni et al., 2019)) without further training on these new datasets.
 (see more details about training on multiple datasets in Appendix D.2.1).
-Due to the lack of zero-shot recommenders in general CF, we slightly modify the two zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022) and UniSRec (Hou et al., 2022), as baselines.
+Due to the lack of zero-shot recommenders in general CF, we slightly modify three zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022), UniSRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), as baselines.
 We also incorporate two strategy-based CF methods (i.e., Random and Pop) and one method using the large language model (LLM) as zero-shot recommender (i.e., LLMRank (Hou et al., 2024b)) (see more details about baselines in Appendix D.2.2).
@@
 UniSRec (Hou et al., 2022) 0.0453 0.0350 0.0863 0.0578 0.1412 0.7135 0.0396 0.0332 0.2454
+VQ-Rec (Hou et al., 2023) 0.0645 0.0410 0.0963 0.0804 0.1921 0.8167 0.0485 0.0492 0.2825
 AlphaRec 0.0913* 0.0573 0.1277* 0.1486* 0.3215* 0.9296* 0.0660* 0.0545* 0.3381*
@@
 As shown in Table 5, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention).
-Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020) by a large margin, even without additional training on search tasks.
-We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more case study results in Appendix D.3.3).
-Additionally, the intention-aware ability also benefits from user historical interests.
-Figure 4(b) depicts the effect of $\alpha$.
-The $\alpha$ controls the strength of user intention, where $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored.
-The convex curve in Figure 4(b) suggests that both user historical interests and user intention play vital roles.
-The above findings suggest the potential of adopting advanced language representations to perceive text-based user intentions and refining recommendations (see more details in Appendix D.3).
+Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020), even without additional training on search tasks.
+We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more examples in Appendix D.3.3).
+Additionally, the intention-aware ability benefits from user historical interests.
+Figure 4(b) depicts the effect of $\alpha$, which controls the strength of user intention.
+Here $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored.
+The convex curve in Figure 4(b) suggests that both user interests and user intention play vital roles.
+Above findings highlight the potential of adopting advanced language representations to perceive text-based user intentions and refine recommendations (see more details in Appendix D.3).
 Table 5: The performance comparison in user intention capture.
@@
 There are several limitations unaddressed in this paper.
-On the one hand, while we explore the relationship between language spaces and behavior spaces through empirical results, there is no theoretical guarantee.
-On the other hand, although we investigate the potential of advanced language representations for recommendation, we do not design any new components for CF models.
+First, this research lacks theoretical guarantee, and we do not contribute to designing any new components for CF models.
+Second, AlphaRec lacks a personalized design, since there is only a single MLP introduced for all users.
+The fixed user intention in Section 5.3 is less practical, lacking exploration on personalized user intentions.
 7 Conclusion
@@
 Inspired by this finding, we discussed how to unleash the power of advanced language representations by developing a simple yet effective CF model called AlphaRec.
-Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: fast convergence, zero-shot ability, and intention-aware ability.
+Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: good initialization for item representations, zero-shot ability, and intention-aware ability.
 Possible future work will involve exploring the space relationship from both theoretical and multimodal perspectives.
 We believed that this paper sheds light on rethinking the connection between language modeling and user behavior modeling, benefiting both natural language processing and recommender system communities.
+Acknowledgments
+This research is supported by the National Science and Technology Major Project (2023ZD0121102), the NExT Research Center, National Natural Science Foundation of China (92270114).
 Appendix A Related Works
@@
 The main idea of linear methods is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure (Vulic et al., 2020) ) from the representations in the hidden layers of LMs, or transforming language representations into another feature space with a linear matrix.
-A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023).
+A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023; Fang et al., 2025).
 Recent studies empirically demonstrate that concepts such as color (Patel & Pavlick, 2022), game states (Li et al., 2023a).
@@
 Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern (Rendle et al., 2012).
-Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a).
-These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024).
+Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a; Zhang et al., 2023c).
+These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024; Zhang et al., 2024a).
 However, although effective, these methods are still limited, due to the ID-based paradigm.
@@
 The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LM-as recommender.
-The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024c).
+The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024d; Liu et al., 2024a; d).
 Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences.
@@
 However, such naive methods tend to yield poor performance compared to traditional models.
-Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023d; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024b; Hua et al., 2023; Chen et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).
+Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023e; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024c; Hua et al., 2023; Chen et al., 2024; Ding et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023f; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).
 Appendix B Uncovering Collaborative Signals in LMs via Linear Mapping
@@
 Especially, for the commercial closed-source model (e.g., text-embedding-ada-v2 and text-embeddings-3-large  [^1] https://platform.openai.com/docs/guides/embeddings (Neelakantan et al., 2022)), we directly call the API interface to obtain representations.
+B.3 Connection with Fine-tuned LLM-based Approaches for Exploring the Relationship Between Language Space and Behavior Space.
+In exploring the space relationship between language modeling and user behavior modeling, there is another research line that uses LLMs to interpret behavior representations (e.g., item representations trained from MF (Koren et al., 2009)) from traditional recommenders (Tennenholtz et al., 2024; Yang et al., 2023; Lei et al., 2024).
+They tend to project behavior representations into the token space of LLMs, and fine-tune LLMs to generate explanations for behavior representations (e.g., generate a movie description given a movie representation trained from MF).
+The fact that LLMs can interpret such behavior representations indicates that the behavior space and language space can be aligned at the token space level, which is similar to the findings in our paper.
+Our paper and these studies can be viewed as complementary works on a similar research goal (i.e., exploration on the connection between language modeling and user behavior modeling from the representation space perspective), providing support for each other.
 Appendix C Leveraging Language Representations for Better Recommendation
@@
 Density 0.0008 0.0026 0.0007 0.0010 0.0542 0.0076
-We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing).
+We incorporate six datasets in this paper, including four datasets from the Amazon platform 222www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing).
 Table 6 reports the dataset statistics.
@@
 These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec).
+We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023).
 For these LM-enhanced CF methods, we adopt the leading method XSimGCL as the backbone.
+We use the same item and user representations as we adopt in AlphaRec for KAR and XSimGCL, to make the comparison relatively fair.
+To evaluate the impact of the modifications, we demonstrate one ablation study on RLMRec in Appendix C.5.
   • MF (Koren et al., 2009; Rendle et al., 2012) is the most basic CF model.
@@
 As illustrated in Figure 6, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.
-C.4 The t-SNE Visualization Comparison
+Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively.
+Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance.
+C.4 Effect of Varying Language Representations in AlphaRec
+We evaluate the effect of different language representations in AlphaRec by varying adopted LMs.
+As reported in Table 7, advanced language representations consistently yield higher performance than early BERT-style language representations.
+Moreover, due to the non-linearity and neighborhood aggregation introduced in AlphaRec, the performance gap between different language representations is narrowed.
+Table 7: Effect of varying language representations in AlphaRec.
+Movies & TV Video Games Books
+Recall NDCG HR Recall NDCG HR Recall NDCG HR
+AlphaRec BERT 0.0994 0.0923 0.4873 0.0960 0.0550 0.2179 0.0719 0.0607 0.3434
+RoBERTa 0.0967 0.0895 0.4793 0.0947 0.0545 0.2167 0.0710 0.0596 0.3386
+Llama2-7B 0.1160 0.1092 0.5388 0.1395 0.0817 0.3003 0.0940 0.0793 0.4081
+Mistral-7B 0.1161 0.1097 0.5421 0.1413 0.0828 0.3020 0.0945 0.0799 0.4090
+text-embedding-ada-v2 0.1152 0.1083 0.5382 0.1437 0.0844 0.3062 0.0933 0.0784 0.4061
+text-embeddings-3-large 0.1221 0.1144 0.5587 0.1519 0.0894 0.3207 0.0991 0.0828 0.4185
+SFR-Embedding-Mistral 0.1225 0.1139 0.5571 0.1521 0.0887 0.3209 0.0982 0.0820 0.4161
+C.5 Ablation Study on the Modification of RLMRec
+We have modified RLMRec (Ren et al., 2024b) by using the same item and user representations as in AlphaRec to make the comparison more fair, rather than using representations of LLM-generated item and user profiles.
+To evaluate the effectiveness of this modification, we conduct an ablation study on the adopted item and user representations.
+Specifically, we implement another version of RLMRec (Ren et al., 2024b) by using the original prompting approach to generate item and user profile, and name this version as RLMRec (LLM-generated profile).
+As shown in Table 8, the performance of RLMRec (LLM-generated profile) is lower than the implementation in this paper.
+Furthermore, the performance of incorporating LLM-generated profiles is not always stable, which may even lead to performance degradation in some cases (e.g., Movies & TV).
+We attribute the performance gap to the possible noise and hallucination introduced by these generated profiles.
+Table 8: Performance comparison of different versions of RLMRec .
+Movies & TV Video Games Books
+Recall NDCG HR Recall NDCG HR Recall NDCG HR
+XSimGCL 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550 0.0879 0.0745 0.3918
+RLMRec (LLM-generated profile) 0.1046 0.0942 0.5063 0.1218 0.0696 0.2692 0.0905 0.0741 0.4049
+RLMRec (This paper) 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997 0.0928 0.0774 0.4092
+C.6 Performance on Amazon Electronic 2023
+To alleviate the information leakage problem that the early datasets may be used in the pretraining data of LMs, we consider the latest version of Amazon 2023 (Hou et al., 2024a) as the newly published dataset.
+This dataset was published in March 2024, which is the latest public dataset that we have access to.
+To further alleviate the information leakage problem, we only consider the interactions after 2022.
+We conduct the same experiment as in Section 4 on this new dataset, and present the performance in Table 9.
+As shown in this table, in this new dataset, AlphaRec still consistently outperforms baselines.
+We attribute this to the fact that language models understand the user preference behind item title description, rather than naively remembering the training data.
+Therefore, AlphaRec still works fine on the latest dataset.
+Table 9: Performance of AlphaRec on Amazon Electronics 2023
+Amazon Electronics 2023
+Recall NDCG HR
+MF (Rendle et al., 2012) $0.0130$ $0.0089$ $0.0136$
+MultVAE (Liang et al., 2018) $0.0227$ $0.0158$ $0.0237$
+LightGCN (He et al., 2021) $0.0237$ $0.0161$ $0.0248$
+SGL (Wu et al., 2021) 0.0519 0.0250 0.0551
+BC Loss (Zhang et al., 2022) 0.0548 0.0265 0.0585
+XSimGCL (Yu et al., 2024) 0.0534 0.0261 0.0569
+KAR (Xi et al., 2023) 0.0611 0.0283 0.0661
+RLMRec (Ren et al., 2024b) 0.0633 0.0288 0.0674
+AlphaRec 0.0687* 0.0323* 0.0732*
+C.7 The t-SNE Visualization Comparison
 (a) LM representations
@@
 D.2.2 Baselines
-Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify two methods in sequential recommendation, ZESRec (Ding et al., 2021) and UniSRec (Hou et al., 2022) as our baselines.
+Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify three methods in sequential recommendation, ZESRec (Ding et al., 2021), UniSRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023) as our baselines.
 Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
@@
 By pre-training on multiple source datasets, UniSRec can conduct zero-shot recommendation on various datasets in a transductive or inductive paradigm.
+  • VQ-Rec (Hou et al., 2023) learns vector-quantized representations for items, enabling transferring between datasets.
+A codebook is learned from language representations, for looking up item embeddings.
 D.2.3 Comparison with LLMRank
@@
 $\tau$ 0.15 0.15 0.2 0.15
+Table 16: Cost for extracting language representations on Amazon Movie & TV
+per 10,000 items Time Cost Money Cost
+Llama2-7B 5 mins 0 \$
+text-embedding-3-large 40 s 0.17 \$
 We conduct all the experiments in PyTorch with a single NVIDIA RTX A5000 (24G) GPU and a 64 AMD EPYC 7543 32-Core Processor CPU.